### v2

# Attenuation Bias

Abstract
Many core concepts in political science are latent and therefore can only be measured with error. Measurement error in a predictor attenuates slope coefficient estimates in regression, biasing them toward zero. We show that widely used strategies for correcting attenuation bias—including instrumental variables and the method of composition—are themselves biased, sometimes even more than simple regression ignoring the measurement error altogether. We derive appropriate bias correction methods using split‑sample measurement strategies. Our approach is modular and can be easily deployed with additive score, factor, or machine learning models, requiring no joint estimation while yielding consistent slopes under standard assumptions. Simulations and applications—political knowledge, democracy indices, and text‑based sentiment—show stronger relationships after correction, sometimes by 50%. Open‑source software implements the procedure. Results underscore that latent predictors demand tailored error correction; otherwise, conventional practice can exacerbate bias.
Keywords: Latent variables; Measurement error; Attenuation bias; Identification restrictions
Word count: 9,999
Introduction
Many key concepts in political science, such as political knowledge, ideology, and democracy, are fundamentally unobservable and can only be estimated based on sets of observed indicators. A large body of research relies on measures of these latent variables as predictors in regression models to test theories and make empirical claims. For example, studies have employed proxies for political knowledge to explain patterns of political participation [11], used estimates of individuals’ ideology to predict vote choice and other political behaviors [46, 26], and used democracy indices to assess the relationship between democratic institutions and economic growth [1] or the risk of conflict [21].
Because these latent variables must be estimated based on observed indicators, these measurements necessarily will include error. Latent variables also have no inherent scale, so researchers must impose identifying restrictions to pin these measurements down. This, combined with measurement error, produces coefficient estimates that are biased in different ways compared to the standard classical measurement error case with observable variables. While the problems of measurement error and attenuation bias are widely recognized [17, 12, 7, 37, 16], the additional challenges posed by latent predictors are often underemphasized in applied work. Indeed, in an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP as well as Political Analysis) between 2000 and 2023, we find that while a growing number of articles mention latent variables or measurement error, a small and relatively constant number discusses the two in combination (see Appendix VII).
In this paper, we examine the issues associated with using estimates of latent variables as predictors and present methods to mitigate the resulting biases under specific assumptions. We build on previous work, including [2] and [5], while deriving important new results and introducing new tools that account for key features of latent variables and their estimation. We show that two commonly used techniques for adjusting regression estimates with latent predictors can and often do exacerbate attenuation bias, sometimes dramatically so. Most notably, the method of composition [47], which is the most prominent method used in this area, actually makes the attenuation bias due to measurement error worse than it would be if the measurement error had simply been ignored altogether.
Building on split-sample methods [3], we use multiple estimates of latent traits based on disjoint sets of indicators to correct standard OLS and 2SLS estimators, which are each biased in opposing directions in the latent predictor case. The method is modular—compatible with additive/averaging scores, factor models, and machine‑learning approaches—and avoids the challenges of full joint estimation of measurement and outcome models (e.g., [9, 42]).
We illustrate these methods with several applications, showing that correcting estimates for measurement error can sometimes result in substantively large changes, in some cases up to 50% in magnitude. But in other cases, measurement error may be mild enough that these corrections are smaller, perhaps even negligible. We also present theory and simulation results that demonstrate that while the naive OLS estimator (ignoring measurement error in the latent predictor) is biased toward zero and previously used methods for addressing measurement error can result in even more severely biased estimates, our corrected estimator dramatically reduces bias and produces estimates very similar to those obtained using the more cumbersome full joint estimation approach. Finally, we provide open-source software for implementing these methods, helping applied researchers account for measurement error when using latent predictors in practice.
1 Setup: Latent Predictors, Identification, and Notation
Our use of the term latent is meant to indicate not only that we do not have a true measure of the variable, but also that the variable is fundamentally unobservable and lacks an objective scale. For example, the values of variables such as knowledge or ideology can only be measured indirectly by examining indicators such as the number of factual questions a respondent answers correctly (for knowledge) or the number of policy questions a respondent gives the conservative response (for ideology). Of course, we could define the variable of interest in these cases as the proportion of questions answered correctly on a particular test (or the proportion of responses given in the conservative direction), which would mean that we do, in fact, have the variable itself measured. But it is likely that we actually care about some true underlying level of knowledge or of liberal-conservative ideology, and that the responses to these indicators are useful because they give us (imperfect) information about each respondent’s true value on these latent scales.
To put it more formally, we use the term latent variable to refer to a variable that is not directly observable, can only be estimated based on a set of stochastic indicators, and that does not have an objectively defined scale, so therefore must be measured on a scale that is defined by one or more identifying restrictions. Using these sorts of variables as predictors in regression models presents additional complications above and beyond the well-known issues with measurement error in (non-latent) predictors, as we discuss below. [^1] We do not consider here the important issue of whether a measurement strategy for a latent trait is theoretically related to the concept of interest (e.g., [36]). Instead, we focus on issues that arise conditional on the measurement strategy chosen by researchers.
It should be noted that, while data-rescaling is usually not a source of bias (see preceding reference to the Maximum Likelihood invariance principle), in this case, the dynamics are different because the scaling factors are data- and measurement error-dependent. [^2] We provide a tutorial outlining this intuition. See: GitHub.com/cjerzak/lpme-software/blob/main/tutorial/IntroTutorial.R?raw=true
Below, we briefly outline the additional complications that arise when using a predictor that is measured with error, is latent, and therefore has no objective scale. The key issue is that some scale must be assumed for the latent trait (i.e., we must pin down the variable’s meaning), both for the true unobserved variable $X$ and for the observed but measured with error variable, $\hat{X}$.
A commonly used identification restriction is that the true latent variable $X$ has mean 0 and variance 1, which identifies the scale in the common one-dimensional setting, except for direction (sign flips). [^3] See [40] for a formal discussion of identification in item response models. The direction of the scale is typically specified as well, with a corresponding identifying restriction imposed, e.g., that higher values of a democracy scale mean that a country is more democratic. In our discussion here, we adopt this typical identification restriction, assuming that the true latent variable $X$ has a mean of 0 and a variance of 1 within the sample. This identification restriction has the advantage of making the latent variable’s scale essentially a $z$-score, indicating how many standard deviations above or below the mean any value is, which can aid interpretability. The results below generally apply similarly to other possible identification restrictions, since changing the identification restrictions will simply impose an affine transformation on the estimated latent traits.
We begin by outlining some key definitions and introducing notation. (Where unambiguous, we omit unit subscripts, $i$.)
  • $X$ is the true (unobserved) latent trait, which is assumed to have mean 0 and standard deviation 1 across observations in the sample.
  • $\tilde{X}$ is the un-transformed estimated latent trait, which is assumed to be equal to the true latent trait plus some mean-zero homoskedastic random error, with the error being independent of the true latent trait. Formally, $\tilde{X}=X+U$ where $U\perp\!\!\!\perp X$, $\mathbb{E}[U]=0$ and $\textrm{Var}(U)=\sigma_{U}^{2}$ so that $\textrm{Var}(\tilde{X})=\textrm{Var}(X)+\textrm{Var}(U)=1+\sigma_{U}^{2}$.
  • $\hat{X}$ is the transformed (i.e., identified) set of estimates of the latent trait that is rescaled to have mean zero and standard deviation one within the sample. In other words, $\hat{X}$ is equal to $\tilde{X}$ minus the sample mean of $\tilde{X}$, all divided by the sample standard deviation of $\tilde{X}$.
Our central parameter of interest is the slope coefficient in a linear regression model predicting some dependent variable $Y$ with the latent trait $X$, which we can write as
$Y_{i}=\beta_{0}+X_{i}\beta_{X}+\epsilon_{i}.$ (1)
However, as described above, the fact that $X$ is a latent variable means that not only is it measured with error, but the scale of our measurements of $X$ is different in subtle ways. Figure 1 presents a visualization of how measurement error operates in this case where the true latent trait $X$ is assumed to have mean 0 standard deviation 1, and the observed estimates of the latent trait $\hat{X}$ are also rescaled to have mean 0 and variance 1.
Figure 1: Illustration of how measurement error operates in the classical case (“With Noise”) and in the case with identification restrictions (“With Noise & Normalization”); normalization is needed in order to fix the scale of the latent trait estimation, but that normalization is data- and measurement-error-dependent.
In the standard (non-latent variable) measurement error setup with an observable predictor, we would observe $\tilde{X}$, which is equal to the true $X$ plus measurement error, and could use it as our predictor. Importantly, note that under these standard assumptions $\tilde{X}$ is on the same scale as, but has larger variance than, the true latent trait $X$ since $\tilde{X}$ is equal to $X$ plus an error term that is independent of $X$. For example, if $X$ was a person’s true weight, we might have $\tilde{X}$ that is the reading from a scale that is equal to the true weight plus some random measurement error. In this case, $\tilde{X}$ would still be on the same scale as $X$ and, importantly, the variance of $\tilde{X}$ would be greater than the variance of $X$. In the case of a latent predictor, not only do we not observe the true variable $X$, but we also do not observe $\tilde{X}$. Instead, we must use the identified and transformed $\hat{X}$ as our predictor, which is rescaled to have the same variance as the true $X$. We can think of $\hat{X}$ as formed by taking the untransformed estimates $\tilde{X}$ and subtracting off the sample mean and dividing by the sample standard deviation. Note that we do not typically know what the variance of $\tilde{X}$ is without knowing the variance of the measurement error $\sigma^{2}_{U}$, so we cannot know what the scale of $\tilde{X}$ would be.
2 Why Attenuation Bias Differs with Latent Predictors
We consider a bivariate (for now) linear regression model predicting some dependent variable $Y$ with the true latent predictor $X$ or, alternatively, with $\tilde{X}$ or with $\hat{X}$, under standard linear model assumptions and assuming that the error term in the linear regression is independent of the measurement error $U$. We use the following notation:
  • $\beta_{X}$, our central parameter of interest, is the true slope coefficient on the (true) latent variable $X$ in the bivariate linear regression predicting $Y$
  • $\hat{\beta}_{X}$ is the least squares estimator of $\beta_{X}$ if we were to use the true $X$ (measured without error) as our predictor
  • $\hat{\beta}_{\tilde{X}}$ is the least squares estimator of $\beta_{X}$ if we were to use the un-transformed estimated latent trait $\tilde{X}$ as our predictor
  • $\hat{\beta}_{\hat{X}}$ is the least squares estimator of $\beta_{X}$ if we were to use the transformed (i.e., identified) latent trait estimates $\hat{X}$ as our predictor
We call $\hat{\beta}_{\hat{X}}$ the naive estimator of the slope since it ignores measurement error in the latent predictor and instead treats point estimates $\hat{X}$ as if they were the true values of the predictor. As mentioned above, in the latent predictor case we generally do not observe either the true latent variable $X$ or the un-transformed estimates $\tilde{X}$, so our discussions of $\hat{\beta}_{X}$ and $\hat{\beta}_{\tilde{X}}$ will be for purposes of illustration and comparison to $\hat{\beta}_{\hat{X}}$ and to the corrected estimators we propose below.
If we were able to use the true latent variable $X$ as our predictor, then the standard linear regression assumptions would ensure that our estimator $\beta_{X}$ of the slope coefficient is unbiased. Of course, the fact that the predictor is latent means that we can never observe it directly, and hence we can never measure it without error.
If, by contrast, we were to use the measured with error predictor $\tilde{X}$, which is on the same scale as the true latent variable $X$, as our predictor, then the slope coefficient in a bivariate regression would have the standard measurement error attenuation factor [17, 10]:
$\mathbb{E}[\hat{\beta}_{\tilde{X}}]=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}\;\beta_{X}=\frac{1}{1+\sigma^{2}_{U}}\;\beta_{X}=\lambda\;\beta_{X}$ (2)
since the identifying assumptions for $X$ include its true values having variance 1. Here $\lambda\equiv\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$ is called the attenuation factor. For example, if on the true latent scale, the measurement error variance were equal to the variance of the latent trait itself (i.e., $\sigma_{U}^{2}=1$ then the attenuation factor would be $\frac{1}{1+1}=\frac{1}{2}$ and the estimated coefficient when using $\tilde{X}$ as the predictor would be half of the size of the true coefficient in expectation.
But when $X$ is a latent variable, we cannot observe $\tilde{X}$. Instead, we observe a noisy but identified estimate of $X$ that has been standardized to have mean 0 and variance 1. Thus, we can think of $\hat{X}$ as being equal to $\frac{\tilde{X}}{\sqrt{1+\sigma_{U}^{2}}}$ (ignoring mean adjustment, which affects only the intercept). It follows that the naive estimator using the identified latent trait estimates is
$\mathbb{E}[\hat{\beta}_{\hat{X}}]=\frac{1}{\sqrt{1+\sigma_{U}^{2}}}\beta_{X}$ (3)
meaning that the attenuation in this case is less severe than in the observable case (since $0<\frac{1}{\sqrt{1+\sigma_{U}^{2}}}>\frac{1}{1+\sigma_{U}^{2}}<1$ for $\sigma_{U}^{2}>0$).
Figure 2: Attenuation factor for estimated slope coefficient as a function of measurement error variance for observable predictor measured with error ($Y$ regressed on $\tilde{X}$) and latent predictor measured with error ($Y$ regressed on $\hat{X}$).
Figure 2 shows the attenuation factor for the bivariate regression slope coefficient estimate as a function of the measurement error variance $\sigma_{U}^{2}$ for the cases of observable and latent predictors. In both cases, it is assumed that the true predictor has variance 1 (equivalently, one can think of the variance of the measurement error as being the ratio of the variance of the measurement error to the variance of the true predictor). When there is no measurement error, the attenuation factor is 1 in both cases. As measurement error increases, both attenuation factors drop below 1, meaning that the expected value of the slope coefficient estimate is smaller in magnitude than the true parameter value. In the latent setting, the scale of the measured-with-error predictor is transformed by the identifying restriction to have variance 1—the same variance as the true variable measured without error. This rescaling multiplies the estimated slope coefficient, making it larger in magnitude and thereby mitigating, but not eliminating, attenuation.
It follows from the different attenuation factors for the latent predictor setup we consider here that applying standard corrections for attenuation based on estimates of measurement error will over-adjust the coefficient, resulting in estimates that are inflated, i.e., biased away from zero. In other words, if one corrects for an attenuation factor of $\frac{1}{1+\sigma^{2}_{U}}$ when the true attenuation factor is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ then this would be an over-correction resulting in amplification bias in the estimated slope coefficient.
3 Existing Approaches and Their Limitations
The most common approach to dealing with measurement error in latent predictors has been to simply ignore the problem and use point estimates of the latent trait as the predictor without any further adjustment. In cases where measurement error is minimal, using this naive estimator may be reasonable. But with significant measurement error, slope estimates will be biased toward zero, sometimes dramatically so, when using standard regression estimates.
A large literature in econometrics and other areas has dealt with so-called errors-in-variables modeling, with the most prominent of these techniques being instrumental variables (see [18] for an overview of these approaches). As mentioned above, the method of composition has been prominently used in recent decades in political science and related literature. Finally, it is sometimes possible to use full joint estimation in which the measurement model for the latent trait and the regression predicting $Y$ with $X$ are estimated together in a unified model. This section briefly discusses each of these approaches, including their benefits and drawbacks.
3.1 Instrumental Variables (IV): Over‑correction under Latent Scaling
One common method for correcting for measurement error in a predictor is instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^4] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [37]). In the standard observable predictor setting, if $X$ is measured with error following the setup above and we seek to estimate the slope coefficient on $X$ in a linear regression predicting some other variable $Y$, we can correct for the resulting attenuation bias by leveraging an instrument $Z$ with the following properties: (i) $\textrm{Cov}(X,Z)\neq 0$; (ii) $\textrm{Cov}(Z,U)=0$; (iii) $\textrm{Cov}(Z,\varepsilon)=0$. In other words, a valid instrument in this setting must be correlated with the true predictor $X$ and be uncorrelated with both the measurement error in $X$ and the error term in the regression predicting $Y$ with $X$.
Standard IV adjustment for measurement error, however, will over-correct in the case of latent predictors. This is because the attenuation factor is different with latent predictors as seen in Equation 3 above. We show below that using standard IV can actually increase bias, resulting in coefficients that are too large in magnitude.
3.2 Method of Composition (MOC): Why It Exacerbates Attenuation
Previous work in political science and elsewhere has considered how to appropriately propagate the uncertainty in estimates of latent predictors into estimated slope coefficients in regression models. Much of this work has employed an approach based on the method of composition (MOC) [45, p. 52]. In a prominent article that introduced the specific variant of MOC commonly used in political science and more broadly, [47] estimate countries’ levels of democracy, a latent trait, using Bayesian MCMC; then, for each iteration of the sampler for the latent traits, they estimate a regression predicting the risk of civil war with level of democracy and other factors, where level of democracy is set to the iteration’s sampled value of the latent traits. Many high-impact works have used MOC with the goal of appropriately accounting for the uncertainty in latent predictors [8, 13, 14, 27, 28, 30, see e.g.]. The Varieties of Democracy (V-Dem) project even provides, in addition to point estimates of their democracy measures, a series of posterior samples so that researchers can employ MOC in subsequent models [38]. [^5] Note that V-Dem’s use of MOC to combine individual measures into more general indices and account for the resulting uncertainty [35] does not necessarily suffer from the issues we identify in using MOC to fit second stage regression models with latent predictors. In a recent APSR Letter, [44] criticize previous work on the relationship between public support for democracy and regime survival, arguing that MOC should be used to incorporate the uncertainty in estimates of the latent predictors.
Although it shares a name with a general method of sampling from a distribution, MOC for latent predictors, as advocated by [47] and as commonly used in the subsequent latent traits literature, uses slightly different steps. The general MOC, used to sample from the density for some random variable $y$, first samples $x^{\prime}$ from the density for some other random variable $g(x)$, then samples $y$ from $f(y|x^{\prime})$. As applied to the latent predictor situation with a matrix of indicators $W$ from which we can estimate the latent trait $X$ according to some statistical model, MOC proceeds according to the following steps for a set of draws $t=1,\ldots,T$:
  1. 1.
    Sample $X^{(t)}$ from $p(X|W)$ according to chosen measurement model
  2. 2.
    Fit a regression predicting $Y$ with $X^{(t)}$, noting the coefficient point estimate $\hat{\beta}^{(t)}$ and its accompanying variance-covariance matrix $\hat{V}^{(t)}$
  3. 3.
    Sample $\beta^{(t)}$ from Multivariate Normal distribution with mean vector $\hat{\beta}^{(t)}$ and variance-covariance matrix $\hat{V}^{(t)}$
Proponents of MOC have argued that the samples $\beta^{(t)}$, typically treated as if they were draws from $p(\beta|W,Y)$, provide a set of draws from a distribution for the slope coefficient that appropriately accounts for the uncertainty in the latent predictor $X$.
MOC is typically motivated by the concern that simply using point estimates of the latent trait as a predictor may result in over-confident findings and that the uncertainty in the latent trait estimates should be taken into account in the estimation, with their uncertainty allowed to “propagate” into the regression level of the model. But as discussed above, the primary concern when estimating regression models using a predictor measured with error is attenuation bias, which results in slope coefficient estimates being pulled toward zero on average for the predictor measured with error. Furthermore, in the simple linear regression case, although standard errors can be biased either upwards or downwards, this is counteracted by the coefficient attenuation, with the overall result that $t$-statistics are biased toward zero. [^6] See [39], p. 3 for a derivation of this $t$-statistic bias. This means that overconfidence is not a problem that typically needs remediation in this situation.
MOC and related approaches average over a set of estimates, each using separate draws from the posterior distribution for the latent variable given the indicator data. Each of these individual draws will typically have even more random variation than do the point estimates of the latent traits. To see this, note that the point estimates are themselves equal to the true latent traits plus some measurement error, and the individual draws from the posterior distribution for $X$ given $W$ can be thought of as being randomly distributed around the point estimates. i.e., $X^{(t)}=\hat{X}+\xi$ where $\hat{X}$ is the point estimate (e.g., the posterior mean for $X$ given $W$) and $\xi$ is an additional source of mean-zero randomness relative to just using the point estimates for $X$. MOC thus averages over a series of even further attenuated estimates as compared to using the point estimates of the latent traits themselves. The attenuation factor with MOC is now based on a measurement error variance that is the sum of the measurement error variance from the traditional case in Equation 3 and the variance of the conditional posterior for the latent traits $X$. In other words, applying MOC is analogous to increasing the measurement error from $\sigma^{2}_{u}$ to $\sigma^{2}_{u}+\sigma^{2}_{\xi}$ for each draw of the sampler, then averaging over these more strongly biased values.
Seen in this light, it is unsurprising that many applications of MOC have estimated slope coefficients on latent predictors to be even closer to zero than those from applications using ordinary regression based on the point estimates of the latent trait. In other words, MOC tends to exacerbate, rather than ameliorate, the main problem of attenuation, generally making it worse (sometimes much worse) than it would be if one had simply ignored measurement error altogether.
3.3 Full Joint Estimation: Benefits and Practical Costs
Although MOC is sometimes described as a joint modeling strategy, it is, strictly speaking, not. In the MOC implementation used with latent predictors, one first draws $X^{(t)}\sim p(X\mid W)$ and then samples $(\beta,\sigma_{\varepsilon}^{2})\sim p(\beta,\sigma_{\varepsilon}^{2}\mid Y,X^{(t)})$. Crucially, $Y$ is not conditioned on when sampling $X^{(t)}$, whereas a genuinely joint approach targets
$p(X,\beta,\theta\mid Y,W)\ \propto\ p(Y\mid X,\beta,\sigma_{\varepsilon}^{2})\,p(W\mid X,\theta)\,p(X)\,p(\beta,\sigma_{\varepsilon}^{2},\theta),$
so that the conditional for the latent trait updates with the outcome, $X\sim p(X\mid W,Y,\beta,\theta)$. Put differently, MOC integrates over $p(X\mid W)$ rather than $p(X\mid W,Y)$ and therefore ignores outcome information when updating $X$, which—given identification-induced rescaling—tends to exacerbate attenuation rather than cure it (as shown above).777[47] make this point and note that MOC effectively assumes $p(X\mid Z,Y)=p(X\mid Z)$; in our notation, their $Z$ corresponds to our indicator matrix $W$.
It is sometimes possible to estimate the measurement and outcome models jointly—for example, with Bayesian factor–outcome models or SEM/LISREL. Under correct specification, such joint estimation conditions on $Y$ when learning $X$ and avoids the extra attenuation that arises when treating a noisy, rescaled $\hat{X}$ as fixed. The costs are practical and conceptual: one must posit a full statistical model including $p(W\mid X)$ and $p(Y\mid X)$, handle scale/sign identification, and fit the joint posterior or likelihood (typically via specialized MCMC or SEM software). This can be demanding, sensitive to misspecification, and often infeasible for “black‑box” scores (e.g., PCA, additive indices, or ML/text embeddings) that lack an explicit likelihood (or one that can be tractably deployed via MCMC).
4 Corrected Estimation Using Split Indicators
One approach to dealing with the specific type of attenuation that arises in the latent predictor case is to fit a standard linear regression using the latent trait estimates ($\hat{X}$) as the predictor and then correct the slope estimate based on the attenuation factor from Equation 3 above. To do this, we need a way to estimate $\sigma^{2}_{U}$ since the attenuation factor is determined by this measurement error variance. Importantly, this is measurement error variance on the scale of the true latent trait $X$, not the transformed (standardized) scale for the estimates of the latent trait that we can observe.
One way to obtain information about $\sigma^{2}_{U}$ is by having multiple measures of the latent trait $X$ for one or more observations. If these measures tend to give very similar values for each observation, then $\sigma^{2}_{U}$ is likely to be small, whereas if there are large differences between the measures, $\sigma^{2}_{U}$ is likely to be larger. In the case of estimating latent traits based on multiple indicators, we can typically obtain multiple measures by partitioning the indicators into multiple sets. For example, imagine we had 10 indicators for a latent trait $X$. We could use 5 of the indicators to obtain estimates $\hat{X}_{1}$ and use the other 5 to obtain estimates $\hat{X}_{2}$. If we split the indicators evenly, it may be reasonable to assume that the variance of the measurement errors for each of these two sets of estimates is the same. [^8] For an odd number of indicators, consider all partitions where the two sets differ by one item.
Following the notation above, we will denote two (unobserved) estimates of the latent trait on the same scale as the true latent trait itself as $\tilde{X}_{1}=X+U_{1}$ and $\tilde{X}_{2}=X+U_{2}$ where we assume that measurement errors $U_{1}$ and $U_{2}$ each have mean zero and the same variance $\sigma^{2}_{U}$ and are independent of each other and of $X$. We also define $\hat{X}_{1}$ and $\hat{X}_{2}$ as the corresponding identified estimates, which we can observe if we assume they have a mean of 0 and a variance of 1. To estimate the attenuation factor, we can first recognize that even without knowing $\tilde{X}_{1}$ or $\tilde{X}_{2}$, the correlation between $\tilde{X}_{1}$ and $\tilde{X}_{1}$ must be the same as the correlation between the observable estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ since these are simply affine transformations of each other. We can then show that
$\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ $=\text{Cor}(\tilde{X}_{1},\tilde{X}_{2})=\frac{\text{Cov}(\tilde{X}_{1},\tilde{X}_{2})}{\sqrt{\sigma^{2}_{\tilde{X}_{1}}\sigma^{2}_{\tilde{X}_{2}}}}$ (4)
$=\frac{\text{Cov}(X+U_{1},X+U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U})^{2}}}$
$=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}=\lambda$
where, as in Equation 2 above, $\lambda$ is the attenuation factor in the standard (non-latent) bivariate linear regression case where the predictor is measured with classical error and, with a slight abuse of notation, $\sigma^{2}_{U}\equiv\textrm{Var}{(U_{1})}=\textrm{Var}{(U_{2})}$. This follows from the assumption that $U_{1}$ and $U_{2}$ are independent of each other and of $X$. In our setup, $\sigma^{2}_{X}=1$ by identification so that $\lambda=\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$.
This, along with Equation 3 above, suggests using the square root of the sample correlation between the two replicate measures of $X$ in order to correct for attenuation in the estimated slope coefficient, $\beta_{X}$. [^9] In some situations, this split-half may not be appropriate. For example, if indicators are designed to tap different aspects of a latent variable rather than with internal consistency in mind (e.g., [20]), it may be preferable to consider test-retest reliability. Noting that in our latent variable predictor setup the attenuation factor for the slope coefficient is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ rather than the $\frac{1}{1+\sigma^{2}_{U}}$ from the standard observable variable case, we could estimate the standard linear regression coefficient using $\hat{X}_{1}$, as our predictor to obtain $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}$ and then correct this estimate to obtain
$\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}=\frac{\hat{\beta}_{\text{OLS},\hat{X}_{1}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}.$ (5)
Here and below, we use the asterisk superscript to denote our corrected estimators. Note that we could have also used $\hat{X}_{2}$ as our predictor instead of $\hat{X}_{1}$ since we have no reason to privilege either of the two sets of estimates. [^10] Note that we cannot use $\hat{X}$—the point estimates of the latent trait using all of the indicators—as our predictor here since we have an estimate of the correction factor based on the amount of measurement error in $\hat{X}_{1}$ and $\hat{X}_{2}$, not in $\hat{X}$ which is estimated based on all of the available indicators rather than one partition of them. We can also average over these two estimators using the two sets of latent traits estimates, which has the benefit of reducing variability from the choice of which partition. [^11] Our averaging approach, and in particular the equivalent averaged instrumental variables (IV) approach we describe in the next section, shares commonalities with [19]. This averaging strategy also has the benefit of mitigating concerns about violations of the assumption that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$, since if such a violation results in our correction factor being too small, then using the other split-half set of indicators would produce a correction factor that is too large and vice-versa. As we show in Appendix I, the bias induced when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ seems likely to be quite small, if not negligible, over plausible ranges of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ when the estimator is averaged across these two values.
Finally, because there are typically multiple ways to partition the indicators into two equally sized sets, we can define our corrected estimator $\hat{\beta}^{*}_{\text{OLS}}$ as the median over all such partitions, where we first average the estimators within each partition between $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{2}}$. This helps guard against the influence of one or more outlying estimates for certain partitions of the indicators, producing more stable estimates.
An advantage of this corrected OLS approach with a correlation-based adjustment is its relative simplicity and ease of implementation, requiring only the ability to split the indicators used to estimate the latent variable into two sets, estimate the latent trait separately using each of the sets of indicators, and calculate the correlation between the two estimates. This means that our approach can be used with virtually any method of estimating a latent trait, from simple additive or averaging approaches that do not specify a formal model to parametric models or advanced ML or AI approaches.
As mentioned above, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. In our case, we might use our two sets of estimates for the latent trait, using one to instrument for the other. But as shown in Appendix II, standard IV adjustments for measurement error will over-correct for attenuation, resulting in estimates that are inflated, i.e. biased away from zero by a factor of $\sqrt{1+\sigma^{2}_{U}}$. This suggests using the correlation between the two sets of latent traits estimates to adjust the standard IV estimate. In contrast to the bivariate OLS case considered above, in the IV case we multiply, rather than divide, our slope estimate by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$. We then define our corrected IV estimator $\hat{\beta}^{*}_{\text{IV}}$ by averaging over the IV indicators using each set of estimates in a given partition as the predictor, then taking the median of those averages over all partitions, as we did above for the corrected OLS estimator.
In fact, as shown in Appendix II, the corrected IV and corrected OLS estimators are actually equivalent in the bivariate case. Despite this equivalence, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. If one predictor is measured with error, we may either have other “control” variables or the predictor measured with error may itself be a control variable in which case adjusting for this may be important to prevent bias in the other coefficient(s) of interest. [^12] Note in the latter case we may not care about appropriately adjusting the IV coefficient on the predictor that is measured with error. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this. In both the OLS and IV cases, we employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.
5 Illustration: Political Knowledge Predicting Duty to Vote
We begin with an example using data from the American National Election Study’s 2024 Time Series Study in order to illustrate our methods, including how our corrected estimator compares to the naive estimator and other approaches. Our latent predictor is political knowledge, which is estimated based on a four-question battery including questions on federal spending, the length of Senate terms, and party control of the House and Senate. [^13] For the full text of items, see Appendix IV.. Many latent traits, such as political knowledge, are often estimated with a small number of indicators in practice. Here, we take the simple approach of estimating the political knowledge level of each respondent with the number of items answered correctly, which we then standardize to have mean 0 and variance 1 within the sample. [^14] To simplify the analysis, we drop all observations with missing values on any of the indicators or the dependent variable, leaving 3,059 of the original 3,349 observations. This standardization makes comparison easier when looking at estimates based on different numbers of indicators as well as different estimation methods, and it also matches the identification restriction used in our derivations above. In applications like this, researchers likely care less about the proportion of correct answers on the specific questions used in the survey and care more about the underlying levels of relative political knowledge held by each respondent, which the question responses provide information about. Given this, identifying the measurement scale through standardization seems a natural approach.
This application demonstrates how our corrected estimators can be applied to any method of estimating a latent trait, including sophisticated models (e.g., item response, factor analytic, etc.) or simpler additive or averaging strategies. Our dependent variable is the response to a 7-point scale question about whether each respondent views voting as a duty or a choice, where we orient the scale so that “Very strongly a duty” is 7 and “Very strongly a choice” is 1. We make no claims here about disentangling the causal relationships between these variables, but only intend this example to illustrate the results of different estimation approaches in commonly encountered situations where latent predictors are estimated based on a small number of indicators.
Figure 3: Grey points show uncorrected linear regression slope coefficients when using various numbers of indicators to estimate the latent trait, then predicting the dependent variable of duty to vote with the point estimates for knowledge. Black dots indicate the averages of these coefficient estimates across all individual estimates using the specified number of indicators to estimate the latent trait. Horizontal lines show the uncorrected OLS estimate $\hat{\beta}_{\hat{X}}$, corrected IV point estimate $\hat{\beta}^{*}_{\text{IV}}$ (which is also equivalent to $\hat{\beta}^{*}_{\text{OLS}}$), and the MOC estimate. See Tables A.IV.3 and A.IV.4 for additional results varying NN in this application.
Naive Estimation (Uncorrected OLS)
We start by fitting a standard linear regression predicting vote duty with estimated knowledge. Here, we are using the point estimates of this latent trait as our predictor without making any corrections to our estimates. In addition to using all four indicators for this estimation, we also separately estimate political knowledge using each possible combination of the four indicators to explore the dynamics of basing our estimates of the latent trait on different numbers of indicators. There are $2^{4}-1=15$ such possibilities since we ignore the case where none of the four questions is included.
Each grey point in Figure 3 plots the point estimate of the slope coefficient from a standard linear regression predicting our dependent variable with estimated knowledge, where knowledge is estimated using a given set of between 1 and 4 of the indicators. The black dots show the average of these OLS slope estimates when using a given number of indicators. We observe that as more items are used to estimate political knowledge, the estimated slope coefficient on political knowledge tends to be higher, which is unsurprising given that a measure using more indicators tends to have less measurement error and, consequently, produce less attenuation than one based on fewer indicators.
The estimate obtained by using all four indicators can be thought of as the naive OLS estimate, since if one were to ignore measurement error, one would presumably use all available indicators to estimate the latent trait, then regress the dependent variable on the point estimates for the latent trait without any further adjustment. This strategy estimates the slope to be 0.42, as indicated in the figure by the horizontal dashed line.
Corrected Split‑Indicator Estimator
We can also create two separate measures of knowledge by partitioning the four indicators into two sets, each including two of the knowledge items. For example, we could define $\hat{X}_{1}$ as the standardized number of correct answers that a respondent obtains on the first two knowledge questions and define $\hat{X}_{2}$ as the standardized number of correct answers that a respondent obtains on the third and fourth knowledge questions. The sample correlation between these two estimates in this case would be 0.25, suggesting a relatively large amount of measurement error. We could then use this correlation to correct the slope coefficient estimate.
The horizontal dotted line in Figure 3 shows the value of our corrected estimator $\hat{\beta}_{\text{IV}}^{*}$, which takes the median of the within-partition averages of the corrected IV estimators over all possible partitions of the four items into two sets of two items. As shown above, this estimator is equivalent to the corrected OLS estimator $\hat{\beta}^{*}_{\text{OLS}}$. This corrected estimate is 0.62, which is nearly 50% larger than the baseline (uncorrected) OLS estimate, showing how accounting for measurement error can yield substantively meaningful changes in our estimates. Note also that weak instruments do not appear to be a problem, given the high correlation between the two sets of latent trait estimates. The $F$-statistic from the first stage averaged well over 200 across all possible partitions, with the smallest $F$-statistic being 134.6—far above the commonly cited threshold of 10 [43].
Had we used the standard (uncorrected) IV estimator (not shown in the figure) we would have estimated the slope as 1.27 on average across these partitions of the indicators, which is more than double the value of corrected estimate. [^15] If we average the uncorrected IV estimates within each partition and then take the median over these within-partition averages, following the strategy for our corrected estimator, the result is 1.14, which is still 85% higher than our corrected estimate. This illustrates the high degree of amplification bias—estimating the slope as being larger in magnitude on average than the true value—that results from applying standard IV-based corrections for measurement error to the latent predictor setting.
Item Response Modeling and Method of Composition
We could also have used an item response model to estimate political knowledge in this application instead of the simple additive strategy. This would entail obtaining point estimates (e.g., maximum likelihood parameter values or posterior means) of the political knowledge levels for each respondent, then using those as predictors in either baseline OLS regressions or IV regressions, as above, to which we could apply appropriate corrections.
Using an item response model to estimate our latent traits here produces baseline OLS slope estimate of 0.45 (quite similar to the 0.42 using the standardized proportion correct over the set of 4 items). [^16] Here we use the ideal function in the pscl R package to estimate latent traits. Using our corrected estimator with the item response model estimates, we estimate the slope as 0.67, which is slightly higher than the corrected estimate using an additive measure rather than the item response model. Estimating an item response model rather than using simple aggregation methods, such as the total number correct, may be expected to provide a more precise estimate of the latent trait, which could account for the slightly larger estimates obtained for both the standard OLS and corrected IV approaches.
When using the latent traits model, we can also apply MOC, which is the primary solution proposed in the political science latent traits modeling literature for accounting for the measurement error inherent in latent predictors. Note that MOC requires some sort of model for estimating the latent trait since we need to be able to obtain draws from the posterior distribution for the latent traits or something similar (e.g., some likelihood-based distribution), so it cannot be applied to simple latent trait estimation strategies like adding or averaging that do not have an explicit model.
But as discussed above, MOC averages over a set of draws from the posterior distribution for the latent traits given the indicators, with each of these draws being an even noisier measure of the true latent trait than the baseline strategy of using the point estimates (e.g., posterior means or maximum likelihood point estimates) for the latent traits as our predictor. In our political knowledge example, MOC yields a point estimate of 0.20, as shown by the solid horizontal line in Figure 3. This is less than half the size of the unadjusted OLS approach, which is itself known to be attenuated, and is less than one-third the size of our corrected estimate. This is consistent with most previous applications of MOC in the literature, which tend to produce coefficient estimates that are smaller in magnitude than the (presumably already attenuated) naive estimates treating latent trait estimates as known, and illustrates how, rather than helping, MOC actually exacerbates the central problem of attenuation bias that arises when using latent predictors in regression models.
This example also suggests that the specific method used to estimate a latent trait (e.g., simple aggregation or averaging versus item response models) may often matter much less for the resulting slope coefficient estimates than does the method of adjusting (or not adjusting) for the inherent measurement error in the latent trait estimates. While there are only minor differences when switching from the simpler aggregation approach to using an item response model here, estimates differed by nearly 50% when moving from naive OLS to our corrected estimator, and the differences were even more dramatic for MOC and unadjusted IV.
6 Simulation: The Role of Sample Size & Dimensionality
In this section, we study how slope estimates for a latent predictor vary with the number of indicators ($M$) and the sample size ($N$), and how different estimation strategies compare. We evaluate (i) naive OLS that treats point estimates of $X$ as known, (ii) uncorrected IV, (iii) the method of composition (MOC), and (iv) our correlation‑adjusted estimator (IV/OLS‑equivalent).
We generate binary indicators using the standard probit ideal‑point model
$\Pr(W_{ij}=1)=\Phi(X_{i}\eta_{j}+\alpha_{j}),$ (6)
and relate the (unobserved) trait to an outcome via
$Y_{i}=\beta_{0}+X_{i}\beta_{X}+\epsilon_{i}.$ (7)
As above, $X$ is latent with no objective scale; we identify it by imposing $\mathbb{E}[X]=0$, $V[X]=1$, and a directional restriction.
To obtain parameter values for our simulations, we first analyze data from the 2016 CCES (N=38,851N{=}38{,}851), which includes 30 agree/disagree policy items that we use as indicators to estimate respondent ideology ($X$). Our dependent variable ($Y$) is presidential vote (0=Democratic, 1=Republican; other responses dropped). The estimates from these analyses are used in these simulations (see Appendix III for more detailed information about this setup and the simulation results). In the largest $N$ and $M$ setting, the naive slope is 0.38, our corrected estimator is 0.40, and the ground‑truth value is 0.40. The small naive–corrected gap in this calibration reflects that M=30M{=}30 yields relatively precise measurement; our simulations below systematically vary $M$ and $N$ to show when corrections matter most.
We fix item parameters $(\eta,\alpha)$ and regression parameters $(\beta_{0},\beta_{X},\sigma^{2}_{\epsilon})$ to their CCES point estimates and vary
$N\in\{10^{2},10^{3},10^{4}\},\qquad M\in\{6,12,30\}.$
For each $(N,M)$, we: (1) sample $N$ respondents and $M$ items without replacement to form $\mathbf{W}\in\{0,1\}^{N\times M}$; (2) estimate item parameters and latent $X$ via EM (binIRT in emIRT) [^17] We use binIRT in emIRT [24] with default priors.; (3) form split‑indicator scores and compute the correlation‑based correction; and (4) estimate $\beta_{X}$ using naive OLS, our corrected estimator $\hat{\beta}^{*}_{\text{IV}}$ (equivalently $\hat{\beta}^{*}_{\text{OLS}}$), uncorrected IV, and MOC. We average results over 16 random partitions of items in the two‑stage procedure.
This setup lets us compare bias and variability across methods as information about $X$ changes along two key axes: more indicators (larger $M$) and more observations (larger $N$). It also highlights when our corrected estimator yields the largest—and substantively meaningful—revisions relative to naive OLS, uncorrected IV, and MOC.
Figure 4 presents the results, focusing on how the absolute bias, standard error of estimation, and RMSE of each of these three estimators vary as the sample size and number of policies change. The standard (uncorrected) OLS estimator shows relatively large absolute bias (note that the actual bias is toward zero, but we use absolute bias to facilitate easier comparison of the magnitude of bias between estimators that are biased in different directions). Recall that the true value for the slope coefficient is 0.40, meaning that the absolute bias around 0.10 as seen for $M=6$ represents a 23.81% change from the true parameter value. This bias does not change meaningfully with the number of observations, but does decrease as the number of indicators used ($M$) grows. Our corrected estimator exhibits very low bias across all values of $M$ and $N$. The standard deviation does not differ meaningfully between the uncorrected OLS and corrected estimates, suggesting that the reduced bias of our proposed estimator is the main difference.
In all cases here, MOC produces worse estimates—notably higher bias and RMSE—than not only our corrected estimator but also naive OLS, which ignores measurement error altogether. This suggests that MOC is not a useful method for correcting for measurement error and, in fact, makes things much worse. Finally, the uncorrected IV estimator has the highest absolute bias across all $M$ and $N$ (as discussed above, the bias of uncorrected IV is away from zero, rather than toward zero as with the naive estimator and MOC). Differences between the standard deviation of the three estimators are relatively small, particularly between the uncorrected OLS and our corrected estimator (the uncorrected IV is much more variable than the other two when $M=6$).
Overall, the RMSE indicates that our corrected estimator outperforms both the other estimators considered across a wide range of $M$ and $N$. Although differences in the RMSE decrease with $M$ (and very slightly with $N$), we see that even with 30 indicators, a value that is higher than the number used in many applications, there are still meaningful relative improvements (e.g., with adjustment decreasing RMSE from 0.025 to <<0.01).
Figure 4: Simulation results varying NN and $M$.
Presenting these simulation results in a different way, Appendix Table A.IV.1 shows the percentage change in coefficient estimates when moving from naive OLS to our corrected estimator, for different values of both $N$ and $M$. While these differences are relatively small when using 30 indicators, we see changes around 20% when $M=12$, 25% when $M=10$, and 35% when $M=6$. Note that estimating latent traits from smaller indicator sets is far more common in practice than using 30 or more indicators. This suggests that accounting for measurement error in latent predictors estimated from binary indicators is likely to make a substantial difference in many applications.
In the Appendix (Figure A.III.3), we also perform various comparisons with other methods not shown in Figure 4, such as the multiple over-imputation and Bayesian joint factor–outcome models requiring simultaneous specification of both measurement and structural components and often depend on intensive MCMC sampling [29, 6]. We find that, in general, overimputation performs well with a large number of factors, but struggles somewhat with smaller $M$, while joint modeling performs best among the group, although it never seems to systematically outperform the adjusted approach outlined here. This suggests that simple partitioning of indicators and attenuation adjustment can be an attractive approach, especially as it does not require complex joint posterior inference, and adjustment factors can be reused for new outcomes.
For a power analysis, see Section A.III.2 of Appendix III. There, we find that the corrected estimator provides improved coverage compared to unadjusted approaches. This improvement does not bring coverage to nominal levels for small values of $M$ (where we undercover, possibly due to small finite-sample bias); for large values of $N$ and $M$, we find nominal coverage for large values of $N$ and $M$ for the corrected estimator, but not others.
7 Further Applications: Democracy and Text Sentiment as Latent Predictors
Up to this point, we have provided some theoretical background, an illustrative example of our proposed method, and generated further evidence through simulation. In the political knowledge application, our proposed correction resulted in very large changes to the estimated slope, suggesting that naive linear regression ignoring measurement error produced a high degree of attenuation bias. In the ideology example that motivated our simulations in Section 6, the corrected estimator produced much smaller changes when the latent trait was estimated using all 30 indicators, but when using fewer indicators to estimate the latent trait, as is more common in the literature, our proposed correction produced substantively large changes.
How significant are these measurement error dynamics in practice when attempting to answer different substantive social science research questions? Here we turn to two more substantive applications—one drawn from comparative politics, another from the realm of text analysis—illustrating how our method can be used to assess whether measurement error in a given application appears to bias estimates to a meaningful degree. Both of these examples also use measurement models for the latent traits that do not have full underlying statistical models. This means that while our corrected estimation approach can be used, other methods, such as MOC or full joint estimation, would not work with these measurement strategies.
7.1 Principal Component Analysis: Democracy and Growth
We here re-analyze the relationship between democracy and growth using this measurement error perspective. Specifically, we investigate whether adjusting for the attenuation bias inherent in latent democracy measures materially alters the estimated impact of democratic institutions on subsequent economic growth. To generate measures of democracy, we use the raw data from the Freedom House Index [22]. The raw indicator-level data collected by index creators, available between 2013 and 2022, provide information on a total of 17 sub-indicators related to electoral processes, political pluralism, the functioning of government, freedom of expression, associational rights, the rule of law, and personal autonomy. Here, we employ a simple measurement strategy, estimating the latent variable of democracy with the first principal component of these indicators. This also provides an illustration of how our approach can be used with measurement strategies for the latent trait of interest that do not have an underlying full statistical model, such as principal components analysis. For the dependent variable, we consider the long difference in log GDP per capita between 2022 and 2013.181818The long difference is: $\Delta\log\bigl(\textrm{GDP per capita}\bigr)\;=\;\log\bigl(\textrm{GDP per capita}_{2022}\bigr)\;-\;\log\bigl(\textrm{GDP per capita}_{2013}\bigr),$ which approximates the cumulative growth rate over the period.
Table 1 shows the naive OLS estimate as well as our corrected estimate of the slope coefficient on democracy. These estimates are nearly identical, with our correction resulting in a half-percentage-point change. In this example, then, our methods suggest that measurement error may be a negligible issue. This is because of the high degree of correlation between the indicators used to measure democracy, which suggests that the latent variable is estimated with considerable precision under the assumptions used here. Appendix Section IV shows that even if we use a smaller subset of the 17 indicators to estimate democracy, the impact of measurement error seems relatively small.
Table 1: Comparison of democracy coefficients (outcome: growth), adjusted and unadjusted. Quantile‐based confidence intervals constructed via bootstrap.
Value Estimate (with CI)
Unadjusted Coefficient 0.0255 [-0.0060, 0.0612]
Adjusted Coefficient 0.0257 [-0.0060, 0.0616]
Change in Estimate 0.520%
7.2 Text Analysis: Sentiment and Incumbent Vote Share
We next analyze an application of text analysis in the realm of political communication. Our primary objective is to determine how sentiment in political rhetoric predicts incumbent success in the next election. Following [15], we might expect a positive coefficient on net positivity in regression analysis; however, negative rhetoric has become increasingly common throughout the 2010s, particularly as social media amplified partisan messaging strategies [25]. We focus the analysis on this period.
For data, we use house‐level general–election returns from the MIT Elections Lab. Congressional speech texts are drawn from the Congressional Record; House member partisanship is drawn from Voteview [33]. After identifying incumbents and computing their district-level two-party vote shares, we aggregate each representative’s House floor speeches by speaker and congressional session. Due to OCR-induced inconsistencies in speaker names preventing full matching, we successfully link on average 350 incumbents’ records per year.
We score each House member’s floor speeches with a pre‑trained RoBERTa sentiment model [34] that outputs probabilities over five labels (‘strongly negative” to ‘strongly positive”). Because many speeches exceed the 512‑token limit, we segment texts, compute segment‑level probabilities, map labels to numeric scores, and average across segments to obtain a speaker‑level sentiment measure. Split samples are formed by repeating this procedure on two non-overlapping halves of the corpora. As this measurement does not rely on a formal statistical model, our corrected estimator is straightforward to implement, whereas MOC, MO, and full joint estimation would require bespoke modifications to large‑scale training pipelines.
Table 2: Comparison of sentiment coefficients, adjusted and unadjusted. Quantile-based confidence intervals constructed via bootstrap.
Value Estimate
Unadjusted Coefficient -2.83 [-4.28, -1.31]
Adjusted Coefficient -3.23 [-4.92, -1.48]
Change in Estimate -14.40%
The naive OLS slope coefficient estimate ignoring measurement error is -2.83. as shown in Table 2. After applying the split-sample correction to adjust for attenuation bias, our corrected estimator is -3.23, which represents a -14.40% increase in magnitude. Here, appropriately accounting for measurement error in the latent predictor yields a small but non-trivial change in our estimates.
8 Guidance, Limitations, Extensions
The derivations and analyses above illustrate the unique challenges encountered when estimating regression models involving a latent variable as a predictor. We demonstrate that using standard regression models with a latent predictor can yield highly attenuated estimates, whereas MOC, which is likely the most prominent method for correction in political science, typically exacerbates these biases. Standard IV corrections can result in estimates that are even more biased in the opposite direction. Our corrected split-sample IV delivers consistent slope estimates, reducing bias (often dramatically) and performing similarly to more cumbersome approaches such as full joint estimation of the latent traits and regression models. In contrast to other approaches, our method can also be used with any estimation strategy for the latent trait, including those without explicit models or likelihoods.
Applied researchers using latent predictors should check standard estimates against our corrected estimates. In some cases, we may expect attenuation bias to be negligible. For example, estimates of legislator ideology based on hundreds of roll call votes tend to be estimated extremely precisely, at least conditional on the assumed measurement models [32]. NOMINATE legislator ideology scores estimated separately with the first and second sessions of the 117th U.S. Senate are correlated at 0.99, implying that naive OLS regressions using NOMINATE scores would be attenuated by less than one percent. Similarly, the example from Section 7.1 above of democracy predicting economic growth shows less than a 1% change in the estimated coefficient after using our corrected estimator.
The estimated coefficients on other latent predictors may show non-trivial, albeit not huge, changes when applying our corrected methods. For example, [5] finds that correlations between various issue scales when estimated from the two halves of ANES items range from 0.61 to 0.84. This implies that naive OLS slope estimates in these situations may need to be adjusted upwards in magnitude by between 9 and 28 percent. As shown in Section 7.2 above, correcting estimates of the relationship between Congressional speech sentiment and subsequent incumbent vote share produces smaller, but still substantively meaningful changes of around 14% in magnitude.
But in other applications for latent predictors, attenuation bias may be much more severe. It is quite common for latent traits to be estimated based on a very small number of indicators. The political knowledge example presented above demonstrates that correcting naive regression estimates using our proposed method yields a roughly 50% increase in the estimated slope coefficient. Furthermore, our simulations in Section 6 demonstrate that when using 10-20 policy questions to estimate respondents’ ideology, our corrected estimators can result in more than 10-25% increases in the magnitude of estimates compared to naive estimation, and when using 6 indicators, correcting estimates can increase the magnitude of estimates by over one third.
There are several avenues for future work. For example, constructing optimal partitions of indicators for our corrected IV approach could minimize the variability of estimates; in settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [31]; these ideas complement our split-indicator strategy. We also plan to explore whether the latent variable case, with its multiple indicators that can be used to construct various sets of IVs, may allow for useful over-identification tests, perhaps of the assumptions necessary for the consistent estimation of the slope coefficient of interest. Finally, we hope to expand our focus on linear models to encompass other regression frameworks (e.g., GLMs) as well as hierarchical models.
Appendix I: Evaluating the Sensitivity of the Corrected Estimator to Violations of $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$
In our main derivations in the paper, particularly in showing how the correlation between the two sets of latent traits estimates can be used to estimate the measurement error variance, we assumed that the measurement error variance was the same for both of our replicate measures $\hat{X}_{1}$ and $\hat{X}_{2}$ which were obtained by partitioning our indicators into two equally sized sets and estimating the latent trait separately based on each of the two sets. Although this may be roughly true given that each set of estimates is based on the same number of indicators (or nearly the same number in the case of an odd number of indicators), it is unlikely that these variances will be exactly the same in practice.
Note, for example, that different indicators may contain different amounts of information about the latent trait. To use an example from educational testing, a question that every single student answers incorrectly does not give us any information about the relative intelligence of the test takers. For example, even if half of test takers answer a question correctly, their responses may not correlate with overall performance—correct answers might reflect regional familiarity with question content or other non-intelligence factors. And, some highly informative indicators may discriminate very strongly between those with low and high levels of the latent trait.
To the extent that the values of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ differ, our estimate of $\lambda$ based on $\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ and hence the implied correction factor applied to the standard OLS or IV estimators will be incorrect which will cause our corrected estimators to be biased. But how large is this bias likely to be in practice if such assumption violations arise? In other words, how different must $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ be to meaningfully affect the ultimate estimates of $\beta_{X}$ that our corrected estimators produce? …and how much are these possible biases mitigated by our strategy of averaging over slope estimates from each of the different partitions of indicators into two sets? In this section, we explore these questions by deriving an expression for the degree to which estimates are affected by given values of the two variances and discussing the implications for our estimators in practice.
Recall that in the main text, we showed that if we estimate a standard linear regression with our observable estimates $\hat{X}$, then we have
$\mathbb{E}[\hat{\beta}_{\hat{X}}$] $=\mathbb{E}[\beta_{\tilde{X}}]\sqrt{1+\sigma^{2}_{U}}\,=\,\frac{1}{1+\sigma^{2}_{U}}\beta_{X}\sqrt{1+\sigma^{2}_{U}}=\frac{1}{\sqrt{1+\sigma^{2}_{U}}}\beta_{X}.$
Since as we show, $\text{Cor}(\hat{X}_{1},\hat{X}_{2})=\frac{1}{1+\sigma^{2}_{U}}$ we can obtain a corrected estimator by dividing $\hat{\beta}_{\hat{X}}$ by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$.
But for the correlation between the two latent traits estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ to appropriately estimate the correction factor in this way, we need to assume that the measurement error variance for both of these estimates of the latent trait are the same, i.e., that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$, which is why we subscript the variance above with $U$ rather than $U_{1}$ or $U_{2}$.
If the variance of the measurement error for $\hat{X}_{1}$ and $\hat{X}_{2}$ is not the same, then the correlation estimates something different. Specifically, we have
$\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ $=\text{Cor}(\tilde{X}_{1},\tilde{X}_{2})=\frac{\text{Cov}(\tilde{X}_{1},\tilde{X}_{2})}{\sqrt{\sigma^{2}_{\tilde{X}_{1}}\sigma^{2}_{\tilde{X}_{2}}}}$ (8)
$=\frac{\text{Cov}(X+U_{1},X+U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\text{Cov}(X,X)+\text{Cov}(X,U_{2})+\text{Cov}(U_{1},X)+\text{Cov}(U_{1},U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\sigma^{2}_{X}+0+0+0}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\sigma^{2}_{X}}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{1}{\sqrt{(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})}},$
since as above we assume by identification that $\sigma^{2}_{X}=1$.
This implies that when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ our proposed correction to the OLS slope estimator will be affected. To see this note that the naive OLS estimator using $\hat{X}_{1}$ as the predictor will be attenuated by a factor of $\frac{1}{\sqrt{1+\sigma^{2}_{U_{1}}}}$ but we if we divide by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ correct the estimator, then we will not cancel out the attenuation factor and instead will be left with
$\frac{\beta}{\sqrt{1+\sigma^{2}_{U_{1}}}}\left[(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})\right]^{\frac{1}{4}}$ $=\beta\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}.$ (9)
The larger the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$, the farther this fraction will be from 1, and hence the more skewed our estimate of this slope coefficient will be.
Figure A.I.1: This figure depicts the value of the factor $\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}$ from Equation 9 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
But how large does the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$ have to be in order to produce a meaningfully large skew from using this correlation-based correction? In order to assess this Figure A.I.1 plots the factor multiplying $\beta$ in 9 as a function of both measurement error variances. The left pane shows the factor multiplying $\beta$ on the right side of 9 as the measurement error variances for the two estimates of the latent trait, each of which ranges between 0 and 1. For most applications, we would expect ex ante that these parameters would fall into this range, i.e., that the measurement error variance for each of the two sets of estimates would not be larger than the sample variance of the true latent trait itself. Within this range, we observe that the factor is small but not always negligible, reaching values below 0.85 (a 15 percent downward skew in the slope estimate) and above 1.15 (a 15 percent upward skew in the slope estimate) for some instances. But note that these values are ones in which one of the measures has measurement error variance very close to 0, and the other has measurement error variance very close to 1. If we were to partition a set of indicators evenly into two sets, it would seem unlikely that we would be in such a situation. But skewing the slope estimate by smaller amounts, e.g., by 5 or 10 percent in either direction, may still be problematic.
The right pane shows the impact of this assumption violation for a much wider (perhaps implausibly so) range of values for the two variances, with each going from 0 to 10. We see that this factor reaches values much farther from 1. Note, however, that for many of these values much farther from 1 we again would be dealing with dramatically differing values for $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$, e.g. 1 and 5, which may be unlikely to occur in most settings.
More importantly, recall that our proposed corrected estimator $\beta_{\hat{X}}^{*}$ averages over the two ways to use a given partition, in the case of the IV formulation, using $\hat{X}_{1}$ to instrument for $\hat{X}_{2}$, and also using $\hat{X}_{2}$ to instrument for $\hat{X}_{1}$. As discussed in the main text, this would be expected to mitigate the skews that result from differing measurement error variances. This is because when these two measurement errors are not equal, the correlation-based correction factor will be too high when using one of the two sets of estimates as the predictor and too low when using the other one. Specifically, when averaging in this way, we would obtain a factor of
$\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}.$ (10)
Figure A.I.2 shows how this new averaged factor varies over values of each variance for correlation-corrected slope estimator that is averaged over using $\hat{X}_{1}$ and using $\hat{X}_{2}$ as the predictor. We see that the factor is much closer to 1 for most plausible parameter values. In the left pane, the factor is between 1 and 1.01 for nearly all possible variances, and is between 1 and 1.002 for a wide range of variances. Even when we let the variances range between 0 and 10, we see that only highly unequal variances produce factors above 1.02. As discussed above, we may be skeptical that measurement error variances would be this unequal in practice when using the same number of indicators to obtain estimates. This suggests that violations of the assumption that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$ are unlikely to be severe enough in practice to have a significant impact on our corrected estimates.
Furthermore, simple and straightforward checks could be used to reassure researchers that a particular partition of indicators into two sets does not result in a large imbalance of information about the latent traits between these two sets of indicators. For example, if using a factor analytic or ideal point model to estimate the latent traits, one could inspect the factor loadings or discrimination parameters from estimating the model using the full set of indicators in order to make sure that one of the two partitions does not contain many that are close to zero and with other containing many that are far from zero. Researchers could also inspect standard errors, posterior standard deviations, or other measures of uncertainty that their chosen measurement model may produce. Even when using simpler strategies, such as additive or averaging, to produce latent trait estimates, inspecting the correlation matrix of the indicators and the distributions of each individual indicator may be enough to prevent the unexpected use of sets of indicators that each contain dramatically different amounts of information about the latent traits. Many of these things are good practice in any application of measurement models, not just for our specific aims of addressing measurement error in latent predictors.
Also note that the expression in Equation 10 holds for a particular partition of the indicators, where the estimates are averaged using one or the other as the predictor. Our proposed corrected estimator $\hat{\beta}^{*}$ takes the median over all such within-partitions averages of the corrected slope estimate, which presumably makes it even less likely that violations of this identical variance assumption would result in large differences in practice.
Overall, then, we can expect that our correlation-corrected estimators (recall that the corrected OLS and IV estimators are equivalent for the bivariate regression case) should perform well in a wide range of situations and that in practical applications, the assumption that these two measurement error variances are identical is unlikely to meaningfully affect our proposed estimators.
Figure A.I.2: This plot displays the value of the factor $\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}$ from Equation 10 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
Appendix II: Corrected Instrumental Variables Estimator
This section illustrates how the standard IV correction for measurement error yields biased upwards in magnitude estimates in the case of latent predictors. It also demonstrates how in the split indicators setup the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ can be used to adjust the IV estimator to provide consistent estimates. Finally, it is shown that the corrected IV estimate is equivalent to the corrected OLS estimate in the bivarate case.
One common method for correcting for measurement error in a predictor is the use of instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^19] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [37]).
In the standard observable predictor setting, if $X$ is measured with error following the setup above and we seek to estimate the slope coefficient on $X$ in a linear regression predicting some other variable $Y$, we can correct for the resulting attenuation bias by leveraging an instrument $Z$ with the following properties: (i) $\textrm{Cov}(X,Z)\neq 0$; (ii) $\textrm{Cov}(Z,U)=0$; (iii) $\textrm{Cov}(Z,\varepsilon)=0$. In other words, a valid instrument in this setting must be correlated with the true predictor $X$ and be uncorrelated with both the measurement error in $X$ and the error term in the regression predicting $Y$ with $X$.
Following [4], the two-stage least squares (2SLS) IV estimator of the slope if we were able to observe $\tilde{X}$ could be written as
$\hat{\beta}_{\textrm{IV},\tilde{X}}=\frac{\hat{\beta}_{Y\sim Z}}{\hat{\beta}_{\tilde{X}\sim Z}}=\frac{\nicefrac{{\textrm{Cov}(Y,Z)}}{{\textrm{Var}(Z)}}}{\nicefrac{{\textrm{Cov}(\tilde{X},Z)}}{{\textrm{Var}(Z)}}}=\frac{\textrm{Cov}(X\beta+\varepsilon,Z)}{\textrm{Cov}(X+U,Z)}=\frac{\beta\textrm{Cov}(X,Z)}{\textrm{Cov}(X,Z)}$ (11)
Since by assumption $\textrm{Cov}(Z,\varepsilon)=0$ and $\textrm{Cov}(Z,U)=0$ and hence it follows that in this observable predictor setup
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\tilde{X}}=\beta_{X}$ (12)
In the latent predictor case, we can obtain a suitable instrumental variable by splitting the indicators used to estimate the latent variable into two sets and estimating the latent trait separately using each of these two sets of indicators, calling these estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ as above. We can then use $\hat{X}_{1}$ as an instrument for $\hat{X}_{2}$ or vice-versa.
Because these are both measures of the same latent trait, they should be highly correlated with each other (which we can also check empirically) and with the true latent trait XX. [^20] In the latent predictor case, different indicators indicators of the same latent trait, and especially scales constructed based on multiple indicators, are typically highly correlated with each other[5, see e.g.], suggesting that the weak instruments problem should rarely be severe in practice. We may also assume that the instrument is independent of the measurement error, i.e., that $\hat{X}_{2}\perp\!\!\!\perp U_{1}\equiv(X-\hat{X}_{1})$.
In contrast to the IV setup with an observable predictor, our estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ are not on the same scale as $X$ itself. If we had $\tilde{X}_{1}=X+U_{1}$ to use as our predictor, then the IV results in Equations 11 and 12 would hold. In our case, however, we use as our predictor $\hat{X}_{1}=\frac{\tilde{X}_{1}}{\sqrt{1+\sigma^{2}_{U}}}$, i.e. a rescaled version of $\tilde{X}$ since it is standardized by dividing by $\sigma_{\tilde{X}}=\sqrt{1+\sigma_{U}^{2}}$. [^21] Note that this is a slight abuse of notation since here $U$ is the measurement error in the latent trait estimates based on the first half of the indicators. In other words, what is being referred to as $\tilde{X}$ (and also $\hat{X}$) has changed to now refer to the latent trait estimates based on half of the indicators. This implies that our IV estimate using $\hat{X}$ will have the following relation with $\hat{\beta}_{\textrm{IV},\tilde{X}}$
$\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\tilde{X}_{1}\mid\tilde{X}_{2}}\sqrt{1+\sigma_{U}^{2}}$ (13)
and hence
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta\sqrt{1+\sigma_{U}^{2}},$ (14)
so that using IV to correct for measurement error in this latent predictor setting actually over-corrects for the attenuation bias in standard linear regression, resulting in coefficient estimates that are inflated in magnitude on average since $\sqrt{1+\sigma_{U}^{2}}>1$ whenever $\sigma_{U}^{2}>0$, i.e., whenever there is any measurement error in the predictor. Note that here, we are using the notation $A\mid B$ to indicate that $A$ is our predictor and $B$ is our instrumental variable.
This implies a further correction that can be accomplished by multiplying the IV estimate by the reciprocal of this factor, so that
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\frac{\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}}{\sqrt{1+\sigma_{U}^{2}}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\sqrt{\lambda}.$ (15)
With this correction in place,
$\operatornamewithlimits{plim}\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta$ (16)
implying that by adjusting the split-half IV estimate to account for the identifying restriction used for the estimated latent traits, which itself is based on the measurement error variance $\sigma^{2}_{U}$, we can obtain a consistent estimate of the linear regression slope coefficient on the true latent $X$ predicting $Y$.
Of course, the correction used in $\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ is based on the variance of the measurement error (on the true latent variable’s scale) $\sigma_{U}^{2}$ which we do not typically know. As discussed in the previous section, however, we use these replicate measures, $\hat{X}_{1}$ and $\hat{X}_{2}$, to learn about $\sigma^{2}_{U}$ and, consequently, about $\lambda$. Since the correlation between these replicate measurements estimates $\lambda$, we can calculate our corrected estimator as the product of the IV slope estimate and the square root of the correlation:
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (17)
Another way of understanding the corrected IV estimator is to note that if we transformed our measures $\hat{X}_{1}$ and $\hat{X}_{2}$ from having variance 1 to having variance $1+\sigma^{2}_{U}$, which would effectively mean we are in the standard measurement error setup, then the standard IV estimator would provide consistent estimates. [^22] We actually would not need to apply this transformation to whichever one is used as the instrument, since any affine transformation of the instrument will cancel out in the expression for the IV estimator. A convenient way to do this is to estimate $\sigma^{2}_{U}$ with $\nicefrac{{1}}{{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}-1]$, then note that since we want the variance of our predictor to be $1+\sigma^{2}_{U}$ we can multiply it by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ and use this transformed predictor in a standard IV analysis. [^23] Since the ratio of the uncorrected OLS and uncorrected IV estimators provides information about $\lambda$ (see [37]), yet another equivalent estimator can be constructed based on the square root of the product of the baseline OLS and uncorrected IV estimators using the partitioned $\hat{X}_{1}$ and $\hat{X}_{2}$. Since this approach is less intuitive, we do not discuss it here. Obviously, multiplying the predictor by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ has the result of dividing the estimated slope coefficient by the same factor, which is equivalent to multiplying it by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ as in Equation 17 above.
As above, we may want to calculate our corrected IV estimator over all equally sized partitions of the indicators for estimating $\hat{X}_{1}$ and $\hat{X}_{2}$. For each such partition take the average of the corrected estimator using $\hat{X}_{1}$ to instrument for $\hat{X}_{2}$ and the one using $\hat{X}_{2}$ to instrument for $\hat{X}_{1}$, then choose as our estimator the median of all such partition averages. In fact, this estimator, which we call $\hat{\beta}^{*}_{\text{IV}}$, is equivalent to $\hat{\beta}_{\text{OLS}}^{*}$ (see Appendix II) In other words, the correlation-adjusted OLS estimator and the correlation-adjusted IV estimator are equivalent and below we refer to them together as the corrected estimator.
Despite the equivalence shown between $\hat{\beta}_{\text{OLS}}^{*}$ and $\hat{\beta}_{\text{IV}}^{*}$ in the bivariate case, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. If one predictor is measured with error, we may either have other “control” variables or the predictor measured with error may itself be a control variable in which case adjusting for this may be important to prevent bias in the other coefficient(s) of interest. [^24] Note in the latter case we may not care about appropriately adjusting the IV coefficient on the predictor that is measured with error. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this. In both the OLS and IV cases, we employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.
Finally, the framework here assumes classical (mean-zero, independent) measurement error. When non-classical error is a concern, identification can still sometimes be recovered under additional conditions; see, for example, [23]. In settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [31]; these ideas complement our split-indicator strategy.
To see that $\hat{\beta}^{*}_{\text{IV}}$, is equivalent to $\hat{\beta}_{\text{OLS}}^{*}$ note that $\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}$, the linear regression slope coefficient estimate when predicting $\hat{X}_{1}$ with $\hat{X}_{2}$, is equal to $\text{Cor}(\hat{X}_{1},\hat{X}_{2})\sqrt{\frac{\text{V}(\hat{X}_{2})}{\text{V}(\hat{X}_{1})}}=\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ since the variance of both $\hat{X}_{1}$ and $\hat{X}_{2}$ are equal to 1 given that each is standardized. This also implies that $\hat{\beta}_{\text{OLS},\hat{X}_{2}\sim\hat{X}_{1}}=\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ since $\text{Cor}(\hat{X}_{1},\hat{X}_{2})=\text{Cor}(\hat{X}_{2},\hat{X}_{1})$ Therefore we can write the uncorrected IV slope coefficient using $\hat{X}_{1}$ as our predictor and $\hat{X}_{2}$ as our instrument as
$\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}}$ (18)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ (19)
and hence we have
$\hat{\beta}^{*}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (20)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}$
$=\hat{\beta}^{*}_{\text{OLS},\hat{X}_{2}}$
Appendix III: Additional Simulation Details & Results
A.III.1.1 Simulation Design Details
We adopt the same data-generating process as above; for each simulation iteration we:
  1. 1.
    Generate simulated matrix of indicators $W^{\textrm{Sim}}$ by randomly drawing from the predictive distribution of the ideal point model in Equation 6.
  2. 2.
    Estimate respondent ideal points $X$ based on the simulated indicators matrix $W^{\textrm{Sim}}$, rescaling these latent traits estimates to have mean 0 and variance 1 with higher values corresponding to more conservative positions, calling these estimates $X^{\textrm{Sim}}$.
  3. 3.
    Estimate the slope coefficient relating $X$ with $Y$ as specified in Equation 1 using simple OLS with the true latent predictor, with the estimated predictor, with the split sample IV (with uncorrected and two corrected estimators).
A.III.1.2 Ideology Estimation
Before discussing a larger-scale simulation that varies the number of observations and factors used in estimation, we first provide a simple illustration of the performance of the estimators discussed above through a set of simulations using a single fixed set of parameter values. To choose these “true” data-generating parameter values, we look at data from the 2016 Cooperative Congressional Election Survey (CCES). This survey includes a large number of policy questions on which respondents are asked their positions (e.g., whether respondents support background checks for all gun sales or whether they think women should always be allowed to have an abortion as a matter of choice). We use 30 of these policy questions as our indicators from which we estimate respondents’ ideologies (ideal points) and the item parameters (difficulty and discrimination parameters) based on the Clinton, Jackman, and Rivers (CJR) ideal point model. [^25] We use the binIRT function in the emIRT package. We drop several policy items that contain a large proportion of missing values. We also drop any respondents having missing values on any of these indicators or on the presidential vote question or who stated they did not vote or voted for someone other than Trump or Clinton, and then, to allow the simulations to run relatively quickly, we randomly select a subset of 1,000 remaining respondents to use. We then use the estimated ideal points to predict 2016 presidential vote (1=Trump, 0=Clinton).
We then set the true values for 1,000 respondents’ ideal points to those point estimates for $X_{i}$ (after imposing the identifying restriction that these ideal points have mean 0, variance 1, and that higher values correspond to more conservative positions) and set the true values of the difficulty ($\eta$) and discrimination parameters ($\alpha$) to those estimated from these indicators in our CCES sample. We also set the true values for the intercept $\beta_{0}$, the slope $\beta_{x}$, and the error variance $\textrm{Var}(\epsilon_{i})$ to roughly correspond to the values estimated in the linear regression predicting presidential vote with the estimated ideal points using the baseline OLS approach. [^26] The estimated slope using our corrected estimates is slightly under 0.4, but for the purposes of the simulations, we use the value of 0.4 since a round number is easier to use as a baseline for comparison in figures and elsewhere. Note that this value for the slope is likely to be attenuated, i.e., smaller in magnitude than the true slope coefficient, since it was estimated from data that itself includes measurement error in the predictor, but we leave it as is for the true value in our simulations.
Although these simulations rely on a single set of true parameters, these values are based on real data of the sort that is often used in latent predictor applications. Furthermore, these simple simulations will illustrate some basic characteristics of the various estimators and their relative performance, at least in this specific example. Figure A.III.1 shows the distributions of simulated slope estimates ($\beta_{X}$) based on 100,000 simulation draws for each estimator.
The top pane of Figure A.III.1 provides a baseline, showing the simulated sampling distribution for the slope coefficient if we actually knew each respondent’s ideal point with certainty (i.e. if there was no measurement error in the predictor) and we estimated a simple bivariate linear regression predicting presidential vote with the ideal points $X$. As expected, the sampling distribution for the slope estimator is centered at the true slope of 0.4. The simulated RMSE is relatively small (0.01). Well over 60 percent of the simulations produced estimates within 0.01 of the true slope, and well over 90 percent produced estimates within 0.02 of the true slope. In a world where the true predictor is measured without error, we can obtain unbiased and relatively precise estimates of the slope coefficient of interest. Of course, with latent predictors, we will never know these true values.
The second pane in Figure A.III.1 plots the simulated distribution of the uncorrected OLS estimator in the case of measurement error. These simulations estimate the slope coefficient on $X$ predicting the presidential vote, assuming we do not actually observe the true latent trait $X$ but instead observe the set of 30 survey questions (indicators) from which we estimate $X$. In other words, we use estimated ideal points $\hat{X}$ as our predictor rather than the true ideal points without any correction or adjustment. As expected, the coefficient estimates are generally smaller in magnitude, with the average estimate falling just below 0.38. The RMSE is more than double that of the known $X$ case.
The third pane of Figure A.III.1 shows estimates from the standard split-sample IV approach without any corrections for the latent variable case. In these simulations, after simulating the matrix of indicators, we split the 30 indicators into the first 15 and the second 15, estimating ideal points for all respondents separately using each of these two sets. This produces $\hat{X_{1}}$ and $\hat{X_{2}}$, two separate measures. We then use two-stage least squares, using $X_{1}$ as an instrument for $X_{2}$ in the model predicting presidential vote. In the standard case of measurement error in an observable (as opposed to latent) predictor, this IV setup can yield consistent estimates of the slope coefficient. But as shown above, when applied to the latent case, it will over-correct for the attenuation bias, producing amplification bias. This is illustrated in the figure, where the average estimate is 0.43; the RMSE is nearly 4 times as large as in the known $X$ case (nearly double the value with uncorrected OLS and estimated $X$).
Finally, the bottom pane of Figure A.III.1 shows split-sample IV estimates corrected using the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ as described in Equation 17. These estimates are quite similar to those using the true $X$, but with a very slightly higher bias and RMSE. This suggests the feasibility of estimating $\beta_{X}$ with the corrected IV approach in the more realistic case where we cannot use the true $X$ to estimate the measurement error variance of the split-half estimates.
In the last two applications (both corrected IV estimates), this simulation suggests favorable bias dynamics. (Appendix I discusses the possibility of a very enduring small amount of downward finite sample bias.)
This example illustrates the attenuation bias induced in estimates of a bivariate linear regression slope by using a noisy measure of a predictor. In our application, the measure is based on a relatively large number (30) of indicators, resulting in a relatively good signal-to-noise ratio. In cases where fewer indicators or less informative indicators are used, the attenuation bias will be larger, as seen below. It is also shown that while the standard IV approach to dealing with measurement error produces over-correction, it is possible to correct for this and obtain slope estimates that are nearly unbiased, with variance barely higher than if the true predictor were known without error.
Figure A.III.1: Illustrative simulation results. The true coefficient is highlighted in green (0.40). The mean of the estimates is highlighted in red across the various approaches.
A.III.1.3 Power Analysis
Figure A.III.2 shows that, across all the simulated settings examined, our corrected procedure retains close to nominal coverage while confidence intervals from the uncorrected OLS and IV estimators frequently fail to contain the true parameter value (0.4) even at moderate sample sizes. With the correct approach, as the number of indicators ($M$) grows, measurement error in the latent predictor diminishes, allowing the corrected estimator’s coverage to stay near nominal levels. In contrast, the uncorrected methods exhibit persistently lower coverage, particularly in smaller $M$ regimes. In this way, statistical inference is improved with the corrections described here.
Figure A.III.2: Power results varying NN and $M$.
A closer look at the small-$M$ regime reveals that even the corrected approach can show mildly reduced coverage there, reflecting finite-sample bias when estimating the latent factor from very little data. One possibility is that, in this low-information setting, random variation in the latent variable estimates via the bootstrap is insufficient to reflect true sampling uncertainty. Another possibility is the presence of small but lingering finite sample bias in this low-information case. Nonetheless, once $N$ and $M$ become sufficiently large, this small-$M$ artifact subsides; the corrected procedure once again not only outperforms the uncorrected OLS or IV, but also yields inferences approaching nominal levels.
A.III.1.4 Comparison with Bayesian Strategies
Researchers have increasingly turned to Bayesian strategies for inference regarding latent traits. We examine three such approaches: (1) method of composition, (2) multiple over-imputation, and (3) joint factor and outcome models. We compare these strategies using the same data-generating process outlined in the simulation section (see Section 6).
The method of composition approach draws from the posterior distribution of the latent trait estimates [47, 45]. We first fit the outcome OLS model to each posterior draw of the latent traits. We then draw random samples from the resulting likelihood distribution, using draws from a Multivariate Gaussian centered at the point estimate and with a heteroskedasticity-consistent variance-covariance matrix. See [47] for details (especially the right panel of p. 215).
Multiple imputation, described in [41] and refined in [9] for the joint outcome and measurement error modeling case, involves generating multiple plausible values for the latent traits using mean and variances from an initial MCMC step as priors—analyzing each imputed dataset separately before combining results.
In the joint modeling strategy, we fit a Bayesian model that simultaneously uses both outcome and observed factor information to generate estimates of the latent quantities of interest. This approach generates a Bayesian posterior over the coefficient estimates. Again, draws of the latent trait, $X$, are rescaled so that the posterior mean takes on a standard deviation of one.
Figure A.III.3: Various approaches, compared against the approaches analyzed in the main text.
Figure A.III.3 presents the results for these approaches. We observe that none of the Bayesian approaches outperform the sample splitting adjustment methods proposed in this paper. These results suggest that Bayesian modeling strategies do not necessarily solve the problems observed in this problem. The Bayesian posterior draws are noisier than the posterior mean estimate, hence increasing the degree of measurement error. Both the joint Bayesian model and multiple imputations face questions around how to perform standardization, whether that be within or across posterior draws.
In contrast, simple OLS estimates based on sample splitting, adjusted for measurement error as proposed in this paper, provide a compelling alternative. Beyond improvements in RMSE, this approach does not require fitting a larger-scale Bayesian model, thereby avoiding the increased computational costs associated with generating uncorrelated draws from the posterior distribution.
Appendix IV: Additional Empirical Results
A.IV.1.5 American Politics: Ideology & Presidential Vote
We here consider an application from American politics, in particular, from the literature on the relationship between latent ideology and partisan vote choice—a relationship that has been extensively studied and shown to be quite strong [26, 46]. To what extent might measurement error in the estimation of latent ideology lead to over- or under-estimation in estimated relationships with the various strategies previously described?
Using the same design as outlined in Section 7.1, we explore the relationship between latent ideology, as measured by 30 questions on the 2016 CCES regarding respondents’ policy positions, and vote choice for president. We again vary the number of observations used and policy factors considered in estimating latent traits, using random sub-samples and computing average estimates. Latent factors are estimated as described above using the expectation-maximization algorithm.
Figure A.IV.1: Measurement–error amplification for partisan‐ideal‐point predictors. Heat maps vary the number of observations (NN, xx‑axis) and the dimensionality of the latent‐space model ($M$, yy‑axis). Left panel: percentage increase (relative to naïve OLS) in the absolute magnitude of the vote‑share coefficient after applying the split‑sample correction; lighter shades indicate larger upward revisions. Right panel: split‑sample correlations between the two halves of the indicator set, with higher values signalling greater measurement stability.
Table A.IV.1: Percentage change in effect estimate of ideology on vote choice, corrected OLS vs. OLS. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 6 10 12 14 18 20 22 24 28 30
$n=$ 80 35.2 23.7 21.3 18.6 14.6 12.7 11.7 10.5 9.0 8.3
225 34.4 24.2 21.2 18.2 14.0 12.6 11.3 9.9 8.2 7.4
632 34.5 25.0 22.0 19.0 14.2 12.5 10.9 9.7 7.8 7.0
1778 34.6 25.4 22.2 19.1 14.1 12.4 10.9 9.5 7.7 6.9
5000 34.4 25.4 22.4 19.2 14.2 12.3 10.8 9.5 7.7 6.9
A.IV.1.6 Comparison: Corrected Estimator and Uncorrected IV in Democracy and Growth Application
We see in Table A.IV.2 that, if using the uncorrected IV strategy, we overestimate the effect of democracy on growth relative to the corrected IV strategy using a split sampling approach. This overestimation is largest when the number of observations and the number of democracy factors used in estimation are small. The underestimation decreases as the number of observations and the number of democracy factors increase, although again at a faster rate for the number of factors than for the number of observations. This again highlights the importance of gathering extensive information around the latent factor of interest.
Figure A.IV.2: Democracy–growth example: Sensitivity to the size of the indicator set. Left panel: percentage change in the estimated effect of democracy on long‑difference GDP growth after the split‑sample correction, plotted against the number of Freedom‑House sub‑indicators used ($M$). Right panel: mean split‑sample correlations for each value of $M$, showing how measurement error shrinks as more indicators are included.
Table A.IV.2: Percentage change in effect estimate of democracy on growth (long difference), corrected IV vs. uncorrected IV. Rows correspond to observation number; columns to dimensions used in estimating latent democratic status.
$M=$ 4 6 8 10 12 14 16
$n=$ 184 -4.8 -3.5 -2.6 -2.1 -1.7 -1.5 -1.3
A.IV.1.7 Effects of Political Knowledge
We here present additional empirical results on the application involving political knowledge.
Table A.IV.3: Percentage change in effect estimate of knowledge on turnout, corrected IV vs. OLS. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 4
$n=$ 80 49.6
225 50.7
632 48.7
1778 49.9
5000 49.8
Table A.IV.4: Percentage change in effect estimate of knowledge on turnout, corrected IV vs. uncorrected IV. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 4
$n=$ 80 -45.0
225 -41.3
632 -41.1
1778 -41.0
5000 -40.8
A.IV.1.8 Additional Information, Knowledge Example
We here provide the full text of questions from the American National Election Study (ANES) data used for the political knowledge and duty to vote application. These data were from the pre-release of the 2024 Time Series Study (downloaded 3/11/2025).
  • V241612
    For how many years is a United States Senator elected - that is, how many years are there in one full term of office for a U.S. Senator? [Respondents specify number of years, 6 is correct answer]
  • V241613
    On which of the following does the U.S. federal government currently spend the least? [Foreign aid (correct), Medicare, National defense, Social Security]
  • V241614
    Do you happen to know which party currently has the most members in the U.S. House of Representatives in Washington? [Democrats, Republicans (correct)]
  • V241615
    Do you happen to know which party currently has the most members in the U.S. Senate? [Democrats (correct), Republicans]
  • V241218x
    V241218x is a summary variable based on V241215, V241216 and V241217.
  • V241215
    Different people feel differently about voting. For some, voting is a duty - they feel they should vote in every election, regardless of their feelings about the candidates and parties. For others, voting is a choice - they feel free to vote or not to vote, depending on their feelings about the candidates and parties. For you personally, is voting mainly a duty, mainly a choice, or neither a duty nor a choice?[Mainly a duty, Mainly a choice, neither a duty nor a choice]
  • V241216
    How strongly do you feel that voting is a duty? [Very strongly, somewhat strongly, or not strongly / Not strongly, somewhat strongly, or very strongly]?
  • V241217
    How strongly do you feel that voting is a choice? [Very strongly, somewhat strongly, or not strongly / Not strongly, somewhat strongly, or very strongly]?
Appendix V: A lpme Package Tutorial
lpme implements a bootstrap analysis for latent variable models with measurement error correction. We average over nPartition split sample partitions:
# install.packages("devtools")
devtools::install_github("cjerzak/lpme/lpme-software")
# Generate data
Yobs <- rnorm(1000)
ObservablesMat <- matrix(sample(c(0, 1), 1000 * 10,
                         replace = TRUE),
                         ncol = 10)
# Latent error correction method, with partitioning and bootstrap
results <- lpme::lpme(
  Yobs,
  ObservablesMat,
  nBoot = 32L,
  nPartition = 10L,
  bootBasis = 1:length(Yobs),
  ReturnIntermediaries = TRUE
# Print the corrected IV coefficient and its standard error
print(c(results\$Corrected_IVRegCoef, results\$Corrected_IVRegSE))
Appendix VI: Code and Data Availability Statement
Additional data and materials will be made available online. For a code implementation, see
for open-source software implementing the methods described in the paper. Replication data and scripts will be made available at a Harvard Dataverse Repository.
Appendix VII: Measurement Error and Latent Predictors in the Political Science Literature
To assess the state of the political science literature regarding measurement error in latent predictors, we conducted an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP) as well as Political Analysis (PA) between 2000 and 2023. We observe in Figure A.VII.1 that the overall percentage of articles mentioning latent variables and, especially, measurement error is increasing, while the number of published works mentioning both measurement error and latent variables has remained relatively constant throughout the study period. This survey of recent political science articles demonstrates that while many researchers are aware of and concerned about measurement error, the full extent of its impact on analyses involving latent variables is, perhaps, not yet widely recognized or addressed in applied work.
Figure A.VII.1: While the overall number of mentions in top political science journals of measurement error and latent variables has been increasing, mentions of the intersection of the two have remained relatively constant in the survey period. The graph shows the number of articles containing each term by year for APSR, AJPS, JOP and PA from 2000 through 2023.