### v3

Abstract
Many core concepts in political science are latent and therefore can only be measured with error. Measurement error in a predictor attenuates slope coefficient estimates in regression, biasing them toward zero. We show that widely used strategies for correcting attenuation bias—including instrumental variables and the method of composition—are themselves biased when applied to latent regressors, sometimes even more than simple regression ignoring the measurement error altogether. We derive a correlation-based correction using split‑sample measurement strategies. Rather than assuming a particular estimation strategy for the latent trait, our approach is modular and can be easily deployed with a wide variety of latent trait measurement strategies including additive score, factor, or machine learning models, requiring no joint estimation while yielding consistent slopes under standard assumptions. Simulations and applications show stronger relationships after our correction, sometimes by as much as 50%. Open‑source software implements the procedure. Results underscore that latent predictors demand tailored error correction; otherwise, conventional practice can exacerbate bias.
Keywords: Latent variables; Measurement error; Attenuation bias; Identification restrictions
Word count: 6,467
Introduction
Many key concepts in political science, such as political knowledge, ideology, and democracy, are fundamentally unobservable and can only be estimated based on sets of observed indicators. A large body of research relies on measures of these latent variables as predictors in regression models to test theories and make empirical claims. For example, studies have employed proxies for political knowledge to explain patterns of political participation [undefk], used estimates of individuals’ ideology to predict vote choice and other political behaviors [undefan, undefv], and used democracy indices to assess the relationship between democratic institutions and economic growth [undef] or the risk of conflict [undeft].
Because latent variables are unobserved and must be inferred from indicators, any empirical proxy for them (e.g., an index score or latent‑trait estimate) inevitably contains measurement error. While the problems of measurement error and attenuation bias are widely recognized [undefp, undefl, undeff, undefae, undefo], the additional challenges posed by latent predictors are often underemphasized in applied work. Indeed, we conducted an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP as well as Political Analysis) between 2000 and 2023. We find that while a growing number of articles mention latent variables or measurement error, a small and relatively constant number discusses the two in combination (see Appendix VII).
In this paper, we examine the bias introduced by using estimated latent variables as predictors and present methods to mitigate it under specific assumptions. We build on previous work, including [undefa] and [undefd], while deriving important new results and introducing new tools that account for key features of latent variables and their estimation. We show that two commonly used techniques for adjusting regression estimates with latent predictors can and often do exacerbate bias, sometimes dramatically so. Most notably, the method of composition [undefao], which is the most prominent method used in this area, typically makes the attenuation bias due to measurement error worse than it would be if the measurement error had simply been ignored altogether.
Building on split-sample methods [undefb], we use multiple estimates of latent traits based on disjoint sets of indicators to correct standard OLS and 2SLS estimators, which are each biased in opposing directions in the latent predictor case. The method is modular—compatible with a wide range of latent trait estimation approaches, including additive/averaging scores, factor models, and machine‑learning approaches—and avoids the challenges of full joint estimation of measurement and outcome models (e.g., [undefh, undefai]).
With this modular correction in hand, we first illustrate its consequences in applications, showing that accounting for latent-predictor measurement error can shift estimated relationships by as much as 50 in magnitude. But in other cases, measurement error may be mild enough that these corrections are smaller, perhaps even negligible. We also present theory and simulation results. These show that with latent predictors the standard OLS estimator (ignoring measurement error in the latent predictor) is biased toward zero, and that previously used methods for addressing measurement error can result in even more severely biased estimates. By contrast, our corrected estimator dramatically reduces bias and produces estimates very similar to those obtained using the more cumbersome full joint estimation approach. Finally, we provide open-source software—an R package, lpmec ( GitHub.com/cjerzak/lpmec-software )—for implementing these methods, helping applied researchers account for measurement error when using latent predictors in practice.
1 Measurement Error and Bias in Linear Regression
1.1 Standard Classical Measurement Error
Before considering the latent predictor case in depth, we briefly review the case of classical measurement error in the standard (non-latent variable) case. We define the standard (non-latent) measurement-error case as where the regressor $X$ has an intrinsic scale but is observed with additive noise $\tilde{X}=X+U$. Here, and throughout, our parameter of interest is the slope $\beta_{X}$ in a linear regression of an outcome $Y$ on a predictor $X$:
$Y_{i}=\beta_{0}+X_{i}\beta_{X}+\epsilon_{i}.$ (1)
We use $\beta_{X}$ to denote the true regression coefficient that we seek to estimate and $X$ is the true predictor variable, measured without error.
Under the the standard classical measurement error setup the true predictor $X$ is not directly observed, but we instead observe
$\tilde{X}=X+U,$ (2)
where $\mathbb{E}[U]=0$, $U$ is independent of both $X$ and $\epsilon$, and $\textrm{V}(U)=\sigma_{U}^{2}$ so that $\textrm{V}(\tilde{X})=\textrm{V}(X)+\textrm{V}(U)$. When estimating a linear regression using noisy $\tilde{X}$, rather than $X$ itself, as our predictor, the least squares estimate of $\beta_{X}$ is biased according to the standard measurement error attenuation factor [undefp, undefj]:
$\operatornamewithlimits{plim}\hat{\beta}_{\tilde{X}}=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}\;\beta_{X}=\lambda\;\beta_{X},$ (3)
where $\lambda\equiv\nicefrac{{\sigma^{2}_{X}}}{{(\sigma^{2}_{X}+\sigma^{2}_{U})}}$ is called the attenuation factor.
1.2 Measurement Error in Latent Predictors
When the predictor $X$ is latent and must be estimated (and then scaled for identification), regression coefficients are attenuated by a different factor than in the standard (observable‑predictor measurement‑error) case. To explore this, we use a simple framework to generalize from the standard measurement error case involving noisy $\tilde{X}$ above.
We remain agnostic about the measurement strategy: $X$ may be reflective (indicators are modeled as responses to $X$) or formative (indicators are combined to define $X$); our results do not depend on the specific strategy used (see [undefi]). Instead, we assume that a researcher already has some approach to estimating the latent variable $X$ based on a set of observed indicators WW. [^1] We do not consider whether a measurement strategy for a latent trait is theoretically related to the concept of interest (e.g., [undefad]). For example, latent variables such as knowledge or ideology can be measured by examining indicators such as the number of factual questions a respondent answers correctly (for knowledge) or the number of policy questions a respondent gives the conservative response (for ideology).
Regardless of the measurement procedure, we model the unstandardized estimate as $\tilde{X}=X+U$, where $U$ captures the estimation error induced by using indicators to infer $X$ rather than observing $X$ directly. We later show that our approach is likely to perform well even when some of the above assumptions about $U$ may be violated.
Importantly, unlike in the standard measurement error case, latent traits variable estimates must be also identified. That is, they must be anchored since they do not have inherent objective scales. We assume that the true latent trait has mean 0 and variance 1 for units in the sample, which identifies the scale in the common one-dimensional setting, except for direction (sign flips). [^2] See [undefag] for a formal discussion of identification in item response models. The direction of the scale is typically specified as well, with a corresponding identifying restriction imposed, e.g., that higher values of a democracy scale mean that a country is more democratic. $\beta_{X}$, our ultimate quantity of interest, is only interpretable relative to the scale of the latent predictor. Multiplying the latent variable by two, for example, would halve its regression coefficient. Therefore, we maintain the simple and interpretable mean 0, standard deviation 1 identification restriction for our proposed method (and for all others examined) to make comparisons more straightforward.
Crucially, this identification restriction applies both to the true latent trait (e.g., imposing $\mathbb{E}[X]=0$ and $\textrm{V}(X)=1$, conceptually) and to whatever point estimates of the latent trait are used in subsequent analyses (e.g., empirically standardizing estimates to have mean zero and variance one). This distinction between the latent scale and the scale of the identified estimates is often ignored in applied work, but it is central to understanding how measurement error operates in the case of latent predictors and to the attenuation factor we derive below.
A natural question is whether this identification-driven rescaling is inevitable; in fact, one could instead report scores on a more “classical” scale—so that their variance matches $\textrm{V}(\tilde{X})=\sigma_{X}^{2}+\sigma_{U}^{2}$—and then apply the usual attenuation correction. Doing so, however, requires knowing (or estimating) $\sigma_{U}^{2}$; and once $\sigma_{U}^{2}$ is available, combining this rescaling with the usual attenuation correction yields, in fact, the same coefficient adjustment as our proposed approach (as we show below).
Formally, we assume that our point estimates of a latent variable predictor are
$\hat{X}=\frac{\tilde{X}-\text{mean}(\tilde{X})}{\text{sd}(\tilde{X})}$, (4)
which are the transformed (i.e., identified) set of estimates of the latent trait. In other words, the estimates have been scaled to have mean zero and standard deviation one within the sample. Consequently, in the latent case the ultimate estimates $\hat{X}$ have the same variance as the true (latent) regressor $X$, whereas in the standard case above, $\textrm{V}(\tilde{X})=\textrm{V}(X)+\textrm{V}(U)>\textrm{V}(X)$. This rescaling changes the attenuation factor in the latent case.
In the latent predictor case, we have $\sigma^{2}_{X}=1$ by identification and since our predictor when estimating our regression will be $\hat{X}$ which is equal to $\tilde{X}$ divided by $\sqrt{1+\sigma^{2}_{U}}$ (ignoring the mean adjustment which only affects the regression intercept), it follows that $\hat{\beta}_{\hat{X}}=\hat{\beta}_{\tilde{X}}\sqrt{1+\sigma^{2}_{U}}$ and thus
$\operatornamewithlimits{plim}\hat{\beta}_{\hat{X}}=\frac{1}{\sqrt{1+\sigma_{U}^{2}}}\beta_{X}.$ (5)
In other words, the slope coefficient estimate in the latent predictor case is still attenuated, but not as severely as in the standard case since $0<\frac{1}{1+\sigma_{U}^{2}}<\frac{1}{\sqrt{1+\sigma_{U}^{2}}}<1$ for $\sigma_{U}^{2}>0$.
In the next section, we review commonly used approaches for handling measurement error with latent predictors and show how several can worsen bias once identification rescaling is taken into account. We then introduce our split‑indicator correlation correction.
2 Existing Approaches and Their Limitations
2.1 No Adjustment
The most common approach to dealing with measurement error in latent predictors has been to simply ignore the problem and use point estimates of the latent trait as the predictor without any further adjustment. In cases where measurement error is small relative to the variance of the predictor, using this standard approach may be reasonable. But with significant measurement error, slope estimates will be biased toward zero, sometimes dramatically so, when using standard regression estimates.
2.2 Full Joint Estimation: Benefits and Practical Costs
When attenuation from treating estimated latent scores as error-free is a concern, the most principled remedy is to estimate the measurement and outcome models jointly—treating $X$ as latent throughout—for example, with Bayesian factor–outcome models [undefy, undefe] or SEM/LISREL. This joint approach is the most principled way to run a regression with a latent predictor. But it is demanding in practice. It requires a full model for the indicators and the outcome, identification choices for scale and sign, and estimation of the joint posterior or likelihood (often via MCMC or SEM). This can be demanding and often infeasible for “black‑box” scores (e.g., PCA, additive indices, or ML/text embeddings) that lack an explicit likelihood (or one that can be tractably deployed via MCMC). As a consequence, researchers often employ other methods.
Because joint estimation is the gold-standard benchmark whenever it is feasible, we treat it as the reference point for evaluating the simpler, two-stage alternatives that we investigate. As we show, our proposed method can provide similar performance (e.g., RMSE) to full joint estimation, while other existing methods perform much worse.
2.3 Instrumental Variables Produce Over-Correction
When full joint estimation is infeasible, a standard errors-in-variables workaround is instrumental variables (IV), which aims to recover $\beta_{X}$ by isolating variation due to the signal rather than measurement noise (cf. [undefq]). In the classical setting with $\tilde{X}=X+U$, IV/2SLS corrects attenuation by using an instrument $Z$ that satisfies: (i) $\textrm{Cov}(X,Z)\neq 0$ (relevance); (ii) $\textrm{Cov}(Z,U)=0$; and (iii) $\textrm{Cov}(Z,\varepsilon)=0$ (exclusion). Under these conditions, 2SLS effectively leverages the part of the observed regressor that co-moves with ZZ—the component attributable to $X$, not UU—to identify the target slope. [^3] Although we focus on measurement error here, IV can also address simultaneity and omitted-variable bias under suitable assumptions (see [undefae]).
This logic is appealing in latent-variable applications because external instruments are often unavailable, but one can sometimes construct internal instruments by producing multiple estimates of the same latent trait from disjoint information. In the classical errors-in-variables case, this “two-measures” strategy can plausibly satisfy the IV conditions: the alternate measure is correlated with $X$ and, if built from independent information, can be approximately uncorrelated with the measurement error in the primary measure (and unrelated to $\varepsilon$).
[…]
This over-correction illustrates a broader theme: classical measurement-error fixes do not automatically carry over to latent predictors once identification rescaling is accounted for. In the next subsection, we show that the method of composition can also worsen bias, but in the opposite direction by exacerbating attenuation. We then introduce our split-indicator correlation correction, which can be viewed as an IV-style strategy that explicitly adjusts for identification-induced rescaling.
2.4 Method of Composition (MOC): Why It Exacerbates Attenuation
A different—and in applied latent-trait work, common—approach is the method of composition (MOC), which aims to propagate uncertainty from the measurement stage, where $X$ is estimated based on a set of observed indicators $W$, into the second-stage regression. In the MOC implementation used with latent predictors, one first draws $X^{(t)}\sim p(X\mid W)$ and then samples $(\beta,\sigma_{\varepsilon}^{2})\sim p(\beta,\sigma_{\varepsilon}^{2}\mid Y,X^{(t)})$. Crucially, $Y$ is not conditioned on when sampling $X^{(t)}$ despite it being informative about $X$ whenever $\beta\neq 0$.444[undefao] note that MOC effectively assumes $p(X\mid W,Y)=p(X\mid W)$ in our notation.
Many high-impact works have used MOC with the goal of appropriately accounting for the uncertainty in latent predictors [undefg, undefm, undefn, undefw, undefx, undefz, undefal, e.g.]. The Varieties of Democracy (V-Dem) project even provides, in addition to point estimates of their democracy measures, a series of posterior samples so that researchers can employ MOC in subsequent analyses [undefaf]. [^5] Our critique concerns using MOC to fit second-stage regression models with latent predictors. V-Dem’s use of MOC to combine individual measures into more general indices and account for the resulting uncertainty [undefac] does not necessarily suffer from the same issues.
As applied to the latent predictor situation with a matrix of indicators $W$ from which we can estimate the latent trait $X$ according to some statistical model, MOC proceeds according to the following steps for a set of draws $t=1,\ldots,T$:
  1. 1.
    Sample $X^{(t)}$ from $p(X|W)$ according to chosen measurement model
  2. 2.
    Fit a regression predicting $Y$ with $X^{(t)}$, noting the coefficient point estimate $\hat{\beta}^{(t)}$ and its accompanying variance-covariance matrix $\hat{V}^{(t)}$
  3. 3.
    Sample $\beta^{(t)}$ from Multivariate Normal distribution with mean vector $\hat{\beta}^{(t)}$ and variance-covariance matrix $\hat{V}^{(t)}$
Proponents of MOC have argued that the samples $\beta^{(t)}$, typically treated as if they were draws from $p(\beta|W,Y)$, provide a set of draws from a distribution for the slope coefficient that appropriately accounts for the uncertainty in the latent predictor $X$.
This two-stage sampling-and-refitting procedure is typically justified as a way to “propagate” measurement-stage uncertainty forward into the regression and avoid overstating precision. However, when the regressor is measured with error, the dominant first‑order problem is attenuation bias: measurement error pulls slope estimates toward zero.
Indeed, MOC averages over a set of estimates, each using separate draws from the posterior distribution for the latent variable given the indicator data. Each of these individual draws will typically have even more random variation than do the point estimates of the latent traits. To see this, note that the point estimates are themselves equal to the true latent traits plus some measurement error, and the individual draws from the posterior distribution for $X$ given $W$ can be thought of as being randomly distributed around the point estimates. i.e., $X^{(t)}=\hat{X}+\xi$ where $\hat{X}$ is the point estimate (e.g., the posterior mean for $X$ given $W$) and $\xi$ is an additional source of mean-zero randomness relative to just using the point estimates for $X$.
3 A Correlation-Corrected Estimator Using Split Indicators
Taken together—naïve IV’s tendency to over-correct once identification rescaling is imposed, and MOC’s tendency to add noise and worsen attenuation—these limitations motivate a different tack: estimate the latent-scale measurement error and then undo the identification-aware attenuation implied by Equation 5. To do this, we need a way to estimate $\sigma^{2}_{U}$, since the attenuation factor is determined by this measurement-error variance. Importantly, this is measurement error variance on the scale of the true latent trait $X$, not the transformed (standardized) scale for the estimates of the latent trait that we can observe.
One way to learn about $\sigma_{U}^{2}$—and therefore about the latent-predictor attenuation factor in Equation 5—is to obtain (at least) two independent measurements of the same latent trait $X$ on the latent scale (i.e., before the identification/standardization step that produces $\hat{X}$). This “two-measures” idea is the same basic ingredient that motivated the IV discussion above. Intuitively, if two measurements of $X$ give very similar values for each observation, then the measurement error variance $\sigma_{U}^{2}$ must be small; if the two measurements often disagree substantially, then $\sigma_{U}^{2}$ must be larger.
In indicator-based latent-trait settings, we can usually construct such repeated measures by partitioning the indicators into two disjoint sets and estimating the trait separately within each set. For example, with 10 indicators for a latent trait $X$, we could use 5 indicators to obtain estimates $\hat{X}_{1}$ and the other 5 to obtain estimates $\hat{X}_{2}$. If we split the indicators evenly, it may be reasonable to assume that the two resulting estimates have the same measurement-error variance. [^6] For an odd number of indicators, consider all partitions where the two sets differ by one item. Following the notation above, let $\tilde{X}_{1}$ and $\tilde{X}_{2}$ denote the two unobserved (unidentified) measurements on the same scale as the true latent trait:
$\tilde{X}_{1}=X+U_{1}\qquad\text{and}\qquad\tilde{X}_{2}=X+U_{2},$
where we assume $U_{1}$ and $U_{2}$ each have mean zero and the same variance $\sigma_{U}^{2}$, and are independent of each other and of XX. [^7] This independence assumption is similar to the conditional independence assumption commonly used in latent traits measurement strategies including many item response models. Finally, let $\hat{X}_{1}$ and $\hat{X}_{2}$ denote the corresponding identified versions of these split-indicator estimates—i.e., the quantities we actually observe after imposing the usual mean-zero, unit-variance normalization (consistent with Equation 4).
Although we do not observe $\tilde{X}_{1}$ or $\tilde{X}_{2}$ directly, their correlation is identical to $\operatorname{Cor}(\hat{X}_{1},\hat{X}_{2})$ because $\hat{X}_{1}$ and $\hat{X}_{2}$ are affine (mean/variance) transformations of $\tilde{X}_{1}$ and $\tilde{X}_{2}$. The idea of then using the correlation between multiple sets of estimates to correct for measurement error has been around since at least [undefaj]. Here we can then show that
$\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ $=\text{Cor}(\tilde{X}_{1},\tilde{X}_{2})=\frac{\text{Cov}(\tilde{X}_{1},\tilde{X}_{2})}{\sqrt{\sigma^{2}_{\tilde{X}_{1}}\sigma^{2}_{\tilde{X}_{2}}}}$ (6)
$=\frac{\text{Cov}(X+U_{1},X+U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U})^{2}}}$
$=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}=\lambda,$
where, as in Equation 3 above, $\lambda$ is the attenuation factor in the standard (non-latent) bivariate linear regression case where the predictor is measured with classical error and, with a slight abuse of notation, $\sigma^{2}_{U}\equiv\textrm{V}{(U_{1})}=\textrm{V}{(U_{2})}$. This follows from the assumption that $U_{1}$ and $U_{2}$ are independent of each other and of $X$. In our setup, $\sigma^{2}_{X}=1$ by identification so that $\lambda=\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$.
In our latent variable predictor setup, the attenuation factor for the slope coefficient is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ (see Equation 5). In the standard observable variable case, it is $\frac{1}{1+\sigma^{2}_{U}}$ (see Equation 3). This suggests using the square root of the sample correlation between the two separate estimates of $X$ to correct attenuation in the estimated slope coefficient, $\beta_{X}$. [^8] In some situations, this split-half may not be appropriate. For example, if indicators are designed to tap different aspects of a latent variable rather than with internal consistency in mind (e.g., [undefs]), it may be preferable to consider test-retest reliability. To do this, we first estimate the standard linear regression coefficient using $\hat{X}_{1}$, as our predictor to obtain $\hat{\beta}_{\hat{X}_{1}}$ and then correct this estimate to obtain
$\hat{\beta}^{*}_{\hat{X}_{1}}=\frac{\hat{\beta}_{\hat{X}_{1}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}.$ (7)
Here and below, we use the asterisk superscript to denote our correlation-corrected estimators. Because we have no reason to privilege either split-half estimate, we compute both $\hat{\beta}^{*}_{\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\hat{X}_{2}}$ and average them within a partition. [^9] We do not use $\hat{X}$—the point estimates of the latent trait using all of the indicators—because our correction factor is estimated from the split-half reliability of $\hat{X}_{1}$ and $\hat{X}_{2}$, not the full-indicator score; applying the same correction to $\hat{X}$ would require additional assumptions or estimation. This averaging reduces variability from the choice of partition. [^10] Our averaging approach, and in particular the equivalent averaged instrumental variables (IV) approach we describe in the next section, shares commonalities with [undefr]. Averaging across the two split-half estimates also makes the method less sensitive to unequal error variances. If one split yields a correction that is too small, the other tends to yield a correction that is too large. As we show in Appendix I, the bias induced when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ seems likely to be quite small, if not negligible, over plausible ranges of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ when the estimator is averaged across these two values.
Finally, because there are typically multiple ways to partition the indicators into two equally sized sets, we can define our corrected estimator $\hat{\beta}^{*}$ as the median over all such partitions, where we first average the estimators within each partition between $\hat{\beta}^{*}_{\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\hat{X}_{2}}$. This helps guard against the influence of one or more outlying estimates for certain partitions of the indicators, producing more stable estimates. We employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.
An advantage of this corrected OLS approach with a correlation-based adjustment is its relative simplicity and ease of implementation. It requires only splitting the indicators used to estimate the latent variable into two sets, estimating the latent trait separately using each set of indicators, and calculating the correlation between the two estimates. This means that our approach can be used with virtually any method of estimating a latent trait, from simple additive or averaging approaches that do not specify a formal model to parametric models or advanced ML or AI approaches. In contrast, a full joint estimation of the latent traits and the regression model will provide the most principled results, but is often quite cumbersome and in cases where the latent trait estimation does not specify a full statistical model this joint approach may not be possible. As we illustrate below, our approach can produce results very similar to full joint estimation.
The same split-indicator idea can also be cast in an IV framework, which is especially useful once we move beyond the bivariate case to multivariate regressions where one or more covariates are measured with error. In our case, we might use our two sets of estimates for the latent trait, using one to instrument for the other. But as shown above standard IV adjustments for measurement error will over-correct for attenuation, resulting in estimates that are inflated, i.e., biased away from zero by a factor of $\sqrt{1+\sigma^{2}_{U}}$. This suggests using the correlation between the two sets of latent trait estimates to adjust the standard IV estimate. In contrast to the bivariate OLS case considered above, in the IV case we multiply, rather than divide, our slope estimate by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$. In fact, as shown in Appendix II, the correlation-corrected IV and correlation-corrected OLS estimators are equivalent in the bivariate case. Despite this equivalence, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. In multivariate regressions, the error-prone variable may be the main predictor or merely a “control”; in either case, failing to account for its measurement error can bias the estimated coefficients on the other covariates of interest. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this.
4 Illustration: Political Knowledge Predicting Duty to Vote
To make the identification and measurement-error issues concrete, we use the 2024 ANES to compare four common approaches in a familiar latent-predictor setting: unadjusted OLS, MOC, uncorrected IV, and our correlation-corrected estimator. The latent predictor is political knowledge, measured with a four-item battery (federal spending, Senate term length, and party control of the House and Senate). [^11] For full item wording, see Appendix IV. As is typical for many latent traits applications, the indicator set is small. We estimate each respondent’s knowledge as the number of correct answers (0–4) and then standardize within-sample to mean 0 and variance 1.121212We drop observations with missing values on any indicator or on the dependent variable, leaving 3,059 of 3,349 respondents. This normalization matches the identifying restriction used above and facilitates comparisons across different indicator subsets and estimation methods. Substantively, we interpret the scale as relative knowledge in the sample, rather than the proportion correct on these specific items.
This application also illustrates the modularity of our correction: it can be applied to any latent-trait estimator, from simple additive scores to item-response or factor models. The outcome $Y$ is a 7-point item on whether voting is a duty or a choice, coded so that “Very strongly a duty” is 7 and “Very strongly a choice” is 1. We make no causal claims; the goal is purely to contrast estimation strategies when a latent predictor is measured with few indicators.
Figure 1: Grey points are uncorrected OLS slope coefficients from regressing duty-to-vote on estimated knowledge, where knowledge is constructed from each nonempty subset of the four indicators (1–4 items). Black dots are the mean slope within each indicator-count. Horizontal lines show: the unadjusted OLS estimate using all four items ($\hat{\beta}_{\hat{X}}$), the correlation-corrected estimate ($\hat{\beta}^{*}$), and the MOC estimate (using an item-response model). See Tables A.IV.3 and A.IV.4 for additional results varying NN.
Baseline: Uncorrected OLS
We begin with the standard practice: regress vote duty on the point estimates of knowledge, without correcting for measurement error. In addition to the all-four-items score, we recompute knowledge for every nonempty subset of the four indicators to show how the slope changes as measurement becomes more (or less) precise. There are $2^{4}-1=15$ such subsets.
Figure 1 shows a clear pattern: as more items are used to construct knowledge, the estimated slope increases. This is consistent with attenuation from measurement error: adding indicators reduces noise and increases the slope toward its latent-scale value. Using all four items (the typical baseline choice) yields an unadjusted OLS slope of 0.42 (dashed line).
Correlation-Corrected Estimator
Next, we form two independent estimates of knowledge by splitting the four items into two pairs, estimating $\hat{X}_{1}$ and $\hat{X}_{2}$ separately, and using their correlation to undo identification-aware attenuation. For example, splitting the first two items versus the last two yields $\operatorname{Cor}(\hat{X}_{1},\hat{X}_{2})=0.25$, indicating substantial measurement error. We compute the correlation-corrected slope within each possible 2–2 split and summarize across splits by taking the median of the within-split averages.
The dotted line in Figure 1 shows the resulting estimate, $\hat{\beta}^{*}=0.62$, nearly 50% larger than the unadjusted OLS estimate. Under the IV interpretation of the split-half procedure, weak instruments do not appear to be a concern: the first-stage $F$-statistic averages well above 200 across partitions, with a minimum of 134.6—far above the conventional threshold of 10 [undefak] (high $F$ here is driven by the large sample size, even with only moderate split-half correlation).
For comparison, applying standard (uncorrected) IV to these split measures (not plotted) produces severe amplification: the average slope across partitions is 1.27, more than double our corrected estimate. [^13] If we instead average uncorrected IV estimates within each partition and then take the median across partitions (mirroring our aggregation strategy), the result is 1.14, still 85% larger than our corrected estimate. This illustrates how conventional IV-style measurement-error corrections can over-correct in the latent-predictor setting.
Item Response Modeling and Method of Composition
Nothing in our correction depends on using an additive score. Alternatively, we can estimate knowledge with an item response model, obtain respondent-level point estimates (e.g., MLEs or posterior means), and then repeat the same baseline and corrected regressions. [^14] For estimation, we use the ideal function from the pscl package in R. With IRT-based scores, the baseline OLS slope is 0.45 (close to 0.42 under the additive score), while the correlation-corrected estimate is 0.67 (slightly above 0.62). These modest shifts are consistent with IRT producing somewhat more precise trait estimates, but the main differences still come from whether measurement error is corrected at all.
With an explicit latent-trait model in hand, we can also apply MOC, the standard recommendation in much of the applied latent-traits literature for propagating uncertainty into second-stage regressions. A practical limitation is that MOC requires posterior draws (or approximations), so it is not available for ad hoc scores like simple sums or averages. More importantly for our purposes, MOC averages over regressions fit to noisier draws of the latent trait than the point estimates, which can exacerbate attenuation. In this application, MOC yields a slope of 0.20 (solid line in Figure 1), less than half the unadjusted OLS estimate and less than one-third of our correlation-corrected estimate. This pattern mirrors many published MOC applications and underscores that, in latent-predictor regressions, MOC can worsen the central first-order problem: attenuation bias from measurement error.
Overall, this example suggests that at least here the choice of latent-trait estimator (simple aggregation versus IRT) matters less for the slope than how measurement error is handled. Moving from unadjusted OLS to our correction shifts the estimated relationship by roughly 50%, while MOC and unadjusted IV push estimates much farther in opposite directions.
5 Comparing the Performance of Estimators with Simulation
To compare the performance of our correlation-corrected estimator with other approaches, in this section we conduct Monte Carlo simulations, generating simulated data ($W$ and $Y$) with parameter values set to point estimates from real data. Specifically, we rely on the 2016 Cooperative Congressional Election Study (CCES) (NN=38,851), a survey that includes 30 agree/disagree policy items that we use as indicators to estimate respondent ideology ($X$). We assume a commonly used item response model
$\Pr(W_{ij}=1)=\Phi(X_{i}\eta_{j}+\alpha_{j}),$ (8)
and our dependent variable ($Y$) in the regression model is reported presidential vote (0=Democratic, 1=Republican; other responses dropped).
Although the type of data and the approach to estimating the latent trait here are quite common in the literature, both the number of observations ($N$) and a number of indicators ($M$) in the CCES are much larger than in most latent traits applications in political science and related literatures. This allows us to assess the performance of estimators across a range of $N$ and $M$ values, from small to large. Specifically, we set parameter values equal to estimates from the full CCES data. We then simulate data for $N\in\{10^{2},10^{3},10^{4}\}$ and $M\in\{6,12,30\}$, sampling observations and indicators randomly for each value of $N$ and $M$. We then estimate $\beta$ with these simulated datasets using each of several approaches discussed above (see Appendix III for more detailed information about this setup and the simulation results).
Figure 2 presents the results, focusing on how absolute bias, standard error, and RMSE vary with sample size $N$ and number of indicators $M$ for each estimator in the comparison (Uncorrected OLS, Cor. Corrected, Method of Composition, and Bayesian, Full). Generally, the standard deviation does not differ much between the estimators, so differences in RMSE are driven primarily by their relative biases. The standard OLS estimator shows relatively large absolute bias (note that the actual bias is toward zero, but we use absolute bias to facilitate easier comparison of the magnitude of bias between estimators that are biased in different directions). The true value of $\beta$ is 0.40, meaning that the absolute bias around 0.10 as seen for $M=6$ represents a 23.81% change relative to the true parameter value. This bias does not change meaningfully with the number of observations, but does decrease as the number of indicators used grows. Our correlation-corrected estimator exhibits very low bias across all values of $M$ and $N$. In fact, the performance of our estimates is very close to that of the full joint estimation strategy, which is often much more cumbersome (or, when the measurement strategy for the latent trait does not specify a full statistical model, may be infeasible). In all cases here, MOC produces worse estimates—notably higher bias and RMSE—than our correlation-corrected estimator. MOC performs worse than standard OLS, which ignores measurement error altogether.
Figure 2: Simulation results varying NN and $M$. Results show comparison of performance of estimators from simluations described in Section 5.
Presenting these simulation results in a different way, Appendix Table A.IV.1 shows the percentage change in coefficient estimates when moving from standard OLS to our corrected estimator, for different values of both $N$ and $M$. While these differences are relatively small when using 30 indicators, we see changes around 20% when $M=12$, 25% when $M=10$, and 35% when $M=6$. Because estimating latent traits from smaller indicator sets is far more common in practice than using 30 or more indicators, these results for $M$ between 6 and 12 are likely to be the empirically relevant regime. This suggests that accounting for measurement error in latent predictors estimated from binary indicators is likely to make a substantial difference in many applications.
In the Appendix (Figure A.III.3), we also perform various comparisons with other methods not shown in Figure 2, such as the multiple over-imputation and standard (uncorrected) IV. We find that, in general, overimputation performs well with a large number of factors, but struggles somewhat with smaller $M$, while uncorrected IV has by far the highest bias (in this case away from zero) and RMSE of any method considered.
For a power analysis, see Section A.III.2 of Appendix III. There, we find that the corrected estimator provides improved coverage compared to unadjusted approaches. This improvement does not bring coverage to nominal levels for small values of $M$ (where we undercover, possibly due to small finite-sample bias); for large values of $N$ and $M$, we find nominal coverage for large values of $N$ and $M$ for our corrected estimator, but not others.
6 Guidance, Limitations, Extensions
The derivations and analyses above illustrate the unique challenges encountered when estimating regression models involving a latent variable as a predictor. We demonstrate that using standard regression models with a latent predictor can yield highly attenuated estimates, whereas MOC, which is likely the most prominent method for correction in political science, typically exacerbates these biases. Standard IV corrections can result in estimates that are even more biased in the opposite direction. Our correlation-corrected estimator reduces bias in slope estimates, often dramatically, and performs similarly to more cumbersome approaches such as full joint estimation of the latent traits and regression models. In contrast to other approaches, our method can also be used with any estimation strategy for the latent trait, including those without explicit models or likelihoods.
Applied researchers using latent predictors should check standard estimates against our corrected estimates. In some cases, we may expect attenuation bias to be negligible. For example, estimates of legislator ideology based on hundreds of roll call votes tend to be estimated extremely precisely, at least conditional on the assumed measurement models [undefab]. NOMINATE legislator ideology scores estimated separately with the first and second sessions of the 117th U.S. Senate are correlated at 0.99, implying that standard OLS regressions using NOMINATE scores would be attenuated by less than one percent. Appendix A.IV.1.6 presents an application of democracy predicting economic growth that shows less than a 1% change in the estimated coefficient after using our corrected estimator.
The estimated coefficients on other latent predictors may show non-trivial, albeit not huge, changes when applying our corrected methods. For example, [undefd] finds that correlations between various issue scales when estimated from the two halves of ANES items range from 0.61 to 0.84. This implies that standard OLS slope estimates in these situations may need to be adjusted upwards in magnitude by between 9 and 28 percent. As shown in Appendix IV, correcting estimates of the relationship between Congressional speech sentiment and subsequent incumbent vote share produces smaller, but still substantively meaningful changes of around 14% in magnitude.
But in other applications for latent predictors, attenuation bias may be much more severe. It is quite common for latent traits to be estimated based on a very small number of indicators. The political knowledge example presented above demonstrates that correcting standard regression estimates using our proposed method yields a roughly 50% increase in the estimated slope coefficient.
Furthermore, our simulations show how correction can matter in practice. With 10–20 policy questions for ideology, the corrected estimates are often 10–25% larger than standard estimates. When using 6 indicators, correcting estimates can increase their magnitude by over one-third.
There are several avenues for future work. One extension is to choose indicator splits that reduce the variance of estimates. Also, when split samples cannot be constructed, researchers can sometimes build instruments from heteroskedasticity. Both ideas fit naturally with our split-indicator approach. We also plan to explore whether the latent variable case, with its multiple indicators that can be used to construct various sets of IVs, may allow for useful over-identification tests, perhaps of the assumptions necessary for the consistent estimation of the slope coefficient of interest. Finally, we hope to expand our focus on linear models to encompass other regression frameworks such as GLMs. $\blacksquare$
Appendix I: Evaluating the Sensitivity of the Corrected Estimator to Violations of $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$
In our main derivations in the paper, particularly in showing how the correlation between the two sets of latent traits estimates can be used to estimate the measurement error variance, we assumed that the measurement error variance was the same for both of our replicate measures $\hat{X}_{1}$ and $\hat{X}_{2}$ which were obtained by partitioning our indicators into two equally sized sets and estimating the latent trait separately based on each of the two sets. Although this may be roughly true given that each set of estimates is based on the same number of indicators (or nearly the same number in the case of an odd number of indicators), it is unlikely that these variances will be exactly the same in practice.
Note, for example, that different indicators may contain different amounts of information about the latent trait. To use an example from educational testing, a question that every single student answers incorrectly does not give us any information about the relative intelligence of the test takers. For example, even if half of test takers answer a question correctly, their responses may not correlate with overall performance—correct answers might reflect regional familiarity with question content or other non-intelligence factors. And, some highly informative indicators may discriminate very strongly between those with low and high levels of the latent trait.
To the extent that the values of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ differ, our estimate of $\lambda$ based on $\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ and hence the implied correction factor applied to the standard OLS or IV estimators will be incorrect which will cause our corrected estimators to be biased. But how large is this bias likely to be in practice if such assumption violations arise? In other words, how different must $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ be to meaningfully affect the ultimate estimates of $\beta_{X}$ that our corrected estimators produce? …and how much are these possible biases mitigated by our strategy of averaging over slope estimates from each of the different partitions of indicators into two sets? In this section, we explore these questions by deriving an expression for the degree to which estimates are affected by given values of the two variances and discussing the implications for our estimators in practice.
Recall that in the main text, we showed that if we estimate a standard linear regression with our observable estimates $\hat{X}$, then we have
$\mathbb{E}[\hat{\beta}_{\hat{X}}$] $=\mathbb{E}[\beta_{\tilde{X}}]\sqrt{1+\sigma^{2}_{U}}\,=\,\frac{1}{1+\sigma^{2}_{U}}\beta_{X}\sqrt{1+\sigma^{2}_{U}}=\frac{1}{\sqrt{1+\sigma^{2}_{U}}}\beta_{X}.$
Since as we show, $\text{Cor}(\hat{X}_{1},\hat{X}_{2})=\frac{1}{1+\sigma^{2}_{U}}$ we can obtain a corrected estimator by dividing $\hat{\beta}_{\hat{X}}$ by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$.
But for the correlation between the two latent traits estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ to appropriately estimate the correction factor in this way, we need to assume that the measurement error variance for both of these estimates of the latent trait are the same, i.e., that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$, which is why we subscript the variance above with $U$ rather than $U_{1}$ or $U_{2}$.
If the variance of the measurement error for $\hat{X}_{1}$ and $\hat{X}_{2}$ is not the same, then the correlation estimates something different. Specifically, we have
$\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ $=\text{Cor}(\tilde{X}_{1},\tilde{X}_{2})=\frac{\text{Cov}(\tilde{X}_{1},\tilde{X}_{2})}{\sqrt{\sigma^{2}_{\tilde{X}_{1}}\sigma^{2}_{\tilde{X}_{2}}}}$ (A.I.1)
$=\frac{\text{Cov}(X+U_{1},X+U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\text{Cov}(X,X)+\text{Cov}(X,U_{2})+\text{Cov}(U_{1},X)+\text{Cov}(U_{1},U_{2})}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\sigma^{2}_{X}+0+0+0}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{\sigma^{2}_{X}}{\sqrt{(\sigma^{2}_{X}+\sigma^{2}_{U_{1}})(\sigma^{2}_{X}+\sigma^{2}_{U_{2}})}}$
$=\frac{1}{\sqrt{(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})}},$
since as above we assume by identification that $\sigma^{2}_{X}=1$.
This implies that when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ our proposed correction to the OLS slope estimator will be affected. To see this note that the naive OLS estimator using $\hat{X}_{1}$ as the predictor will be attenuated by a factor of $\frac{1}{\sqrt{1+\sigma^{2}_{U_{1}}}}$ but we if we divide by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ correct the estimator, then we will not cancel out the attenuation factor and instead will be left with
$\frac{\beta}{\sqrt{1+\sigma^{2}_{U_{1}}}}\left[(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})\right]^{\frac{1}{4}}$ $=\beta\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}.$ (A.I.2)
The larger the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$, the farther this fraction will be from 1, and hence the more skewed our estimate of this slope coefficient will be.
Figure A.I.1: This figure depicts the value of the factor $\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}$ from Equation A.I.2 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
But how large does the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$ have to be in order to produce a meaningfully large skew from using this correlation-based correction? In order to assess this Figure A.I.1 plots the factor multiplying $\beta$ in Equation A.I.2 as a function of both measurement error variances. The left pane shows the factor multiplying $\beta$ on the right side of Equation A.I.2 as the measurement error variances for the two estimates of the latent trait, each of which ranges between 0 and 1. For most applications, we would expect ex ante that these parameters would fall into this range, i.e., that the measurement error variance for each of the two sets of estimates would not be larger than the sample variance of the true latent trait itself. Within this range, we observe that the factor is small but not always negligible, reaching values below 0.85 (a 15 percent downward skew in the slope estimate) and above 1.15 (a 15 percent upward skew in the slope estimate) for some instances. But note that these values are ones in which one of the measures has measurement error variance very close to 0, and the other has measurement error variance very close to 1. If we were to partition a set of indicators evenly into two sets, it would seem unlikely that we would be in such a situation. But skewing the slope estimate by smaller amounts, e.g., by 5 or 10 percent in either direction, may still be problematic.
The right pane shows the impact of this assumption violation for a much wider (perhaps implausibly so) range of values for the two variances, with each going from 0 to 10. We see that this factor reaches values much farther from 1. Note, however, that for many of these values much farther from 1 we again would be dealing with dramatically differing values for $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$, e.g. 1 and 5, which may be unlikely to occur in most settings.
More importantly, recall that our proposed corrected estimator $\beta_{\hat{X}}^{*}$ averages over the two ways to use a given partition, in the case of the IV formulation, using $\hat{X}_{1}$ to instrument for $\hat{X}_{2}$, and also using $\hat{X}_{2}$ to instrument for $\hat{X}_{1}$. As discussed in the main text, this would be expected to mitigate the skews that result from differing measurement error variances. This is because when these two measurement errors are not equal, the correlation-based correction factor will be too high when using one of the two sets of estimates as the predictor and too low when using the other one. Specifically, when averaging in this way, we would obtain a factor of
$\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}.$ (A.I.3)
Figure A.I.2 shows how this new averaged factor varies over values of each variance for correlation-corrected slope estimator that is averaged over using $\hat{X}_{1}$ and using $\hat{X}_{2}$ as the predictor. We see that the factor is much closer to 1 for most plausible parameter values. In the left pane, the factor is between 1 and 1.01 for nearly all possible variances, and is between 1 and 1.002 for a wide range of variances. Even when we let the variances range between 0 and 10, we see that only highly unequal variances produce factors above 1.02. As discussed above, we may be skeptical that measurement error variances would be this unequal in practice when using the same number of indicators to obtain estimates. This suggests that violations of the assumption that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$ are unlikely to be severe enough in practice to have a significant impact on our corrected estimates.
Furthermore, simple and straightforward checks could be used to reassure researchers that a particular partition of indicators into two sets does not result in a large imbalance of information about the latent traits between these two sets of indicators. For example, if using a factor analytic or ideal point model to estimate the latent traits, one could inspect the factor loadings or discrimination parameters from estimating the model using the full set of indicators in order to make sure that one of the two partitions does not contain many that are close to zero and with other containing many that are far from zero. Researchers could also inspect standard errors, posterior standard deviations, or other measures of uncertainty that their chosen measurement model may produce. Even when using simpler strategies, such as additive or averaging, to produce latent trait estimates, inspecting the correlation matrix of the indicators and the distributions of each individual indicator may be enough to prevent the unexpected use of sets of indicators that each contain dramatically different amounts of information about the latent traits. Many of these things are good practice in any application of measurement models, not just for our specific aims of addressing measurement error in latent predictors.
Also note that the expression in Equation A.I.3 holds for a particular partition of the indicators, where the estimates are averaged using one or the other as the predictor. Our proposed corrected estimator $\hat{\beta}^{*}$ takes the median over all such within-partitions averages of the corrected slope estimate, which presumably makes it even less likely that violations of this identical variance assumption would result in large differences in practice.
Overall, then, we can expect that our correlation-corrected estimators (recall that the corrected OLS and IV estimators are equivalent for the bivariate regression case) should perform well in a wide range of situations and that in practical applications, the assumption that these two measurement error variances are identical is unlikely to meaningfully affect our proposed estimators.
Figure A.I.2: This plot displays the value of the factor $\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}$ from Equation A.I.3 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
Appendix II: Corrected Instrumental Variables Estimator
This section illustrates how the standard IV correction for measurement error yields biased upwards in magnitude estimates in the case of latent predictors. It also demonstrates how in the split indicators setup the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ can be used to adjust the IV estimator to provide consistent estimates. Finally, it is shown that the correlation-corrected IV estimate is equivalent to the correlation-corrected OLS estimate in the bivarate case.
One common method for correcting for measurement error in a predictor is the use of instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^15] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [undefae]).
In the standard observable predictor setting, if $X$ is measured with error following the setup above and we seek to estimate the slope coefficient on $X$ in a linear regression predicting some other variable $Y$, we can correct for the resulting attenuation bias by leveraging an instrument $Z$ with the following properties: (i) $\textrm{Cov}(X,Z)\neq 0$; (ii) $\textrm{Cov}(Z,U)=0$; (iii) $\textrm{Cov}(Z,\varepsilon)=0$. In other words, a valid instrument in this setting must be correlated with the true predictor $X$ and be uncorrelated with both the measurement error in $X$ and the error term in the regression predicting $Y$ with $X$.
Following [undefc], the two-stage least squares (2SLS) IV estimator of the slope if we were able to observe $\tilde{X}$ could be written as
$\hat{\beta}_{\textrm{IV},\tilde{X}}=\frac{\hat{\beta}_{Y\sim Z}}{\hat{\beta}_{\tilde{X}\sim Z}}=\frac{\nicefrac{{\textrm{Cov}(Y,Z)}}{{\textrm{V}(Z)}}}{\nicefrac{{\textrm{Cov}(\tilde{X},Z)}}{{\textrm{V}(Z)}}}=\frac{\textrm{Cov}(X\beta+\varepsilon,Z)}{\textrm{Cov}(X+U,Z)}=\frac{\beta\textrm{Cov}(X,Z)}{\textrm{Cov}(X,Z)}.$ (A.II.1)
Since by assumption $\textrm{Cov}(Z,\varepsilon)=0$ and $\textrm{Cov}(Z,U)=0$ and hence it follows that in this observable predictor setup
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\tilde{X}}=\beta_{X}.$ (A.II.2)
In the latent predictor case, we can obtain a suitable instrumental variable by splitting the indicators used to estimate the latent variable into two sets and estimating the latent trait separately using each of these two sets of indicators, calling these estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ as above. We can then use $\hat{X}_{1}$ as an instrument for $\hat{X}_{2}$ or vice-versa.
Because these are both measures of the same latent trait, they should be highly correlated with each other, which we can also check empirically, and also highly correlated with the true latent trait XX. [^16] In the latent predictor case, different indicators indicators of the same latent trait, and especially scales constructed based on multiple indicators, are typically highly correlated with each other [undefd, see e.g.], suggesting that the weak instruments problem should rarely be severe in practice. We may also assume that the instrument is independent of the measurement error, i.e., that $\hat{X}_{2}\perp\!\!\!\perp U_{1}\equiv(X-\hat{X}_{1})$. [^17] This is analogous to the conditional independence assumption used in many latent traits models, i.e. that conditional on $X$ the indicators are independent of each other.
In contrast to the IV setup with an observable predictor, our estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ are not on the same scale as $X$ itself. If we had $\tilde{X}_{1}=X+U_{1}$ to use as our predictor, then the IV results in Equations A.II.1 and A.II.2 would hold. In our case, however, we use as our predictor $\hat{X}_{1}=\frac{\tilde{X}_{1}}{\sqrt{1+\sigma^{2}_{U}}}$, i.e. a rescaled version of $\tilde{X}$ since it is standardized by dividing by $\sigma_{\tilde{X}}=\sqrt{1+\sigma_{U}^{2}}$. [^18] Note that this is a slight abuse of notation since here $U$ is the measurement error in the latent trait estimates based on the first half of the indicators. In other words, what is being referred to as $\tilde{X}$ (and also $\hat{X}$) has changed to now refer to the latent trait estimates based on half of the indicators. This implies that our IV estimate using $\hat{X}$ will have the following relation with $\hat{\beta}_{\textrm{IV},\tilde{X}}$
$\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\tilde{X}_{1}\mid\tilde{X}_{2}}\sqrt{1+\sigma_{U}^{2}}$ (A.II.3)
and hence
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta\sqrt{1+\sigma_{U}^{2}},$ (A.II.4)
so that using IV to correct for measurement error in this latent predictor setting actually over-corrects for the attenuation bias in standard linear regression, resulting in coefficient estimates that are inflated in magnitude on average since $\sqrt{1+\sigma_{U}^{2}}>1$ whenever $\sigma_{U}^{2}>0$, i.e., whenever there is any measurement error in the predictor. Note that here, we are using the notation $A\mid B$ to indicate that $A$ is our predictor and $B$ is our instrumental variable.
This implies a further correction that can be accomplished by multiplying the IV estimate by the reciprocal of this factor, so that
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\frac{\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}}{\sqrt{1+\sigma_{U}^{2}}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\sqrt{\lambda}.$ (A.II.5)
With this correction in place,
$\operatornamewithlimits{plim}\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta,$ (A.II.6)
implying that by adjusting the split-half IV estimate to account for the identifying restriction used for the estimated latent traits, which itself is based on the measurement error variance $\sigma^{2}_{U}$, we can obtain a consistent estimate of the linear regression slope coefficient on the true latent $X$ predicting $Y$.
Of course, the correction used in $\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ is based on the variance of the measurement error (on the true latent variable’s scale) $\sigma_{U}^{2}$ which we do not typically know. As discussed in the previous section, however, we use these replicate measures, $\hat{X}_{1}$ and $\hat{X}_{2}$, to learn about $\sigma^{2}_{U}$ and, consequently, about $\lambda$. Since the correlation between these replicate measurements estimates $\lambda$, we can calculate our corrected estimator as the product of the IV slope estimate and the square root of the correlation:
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (A.II.7)
Another way of understanding the corrected IV estimator is to note that if we transformed our measures $\hat{X}_{1}$ and $\hat{X}_{2}$ from having variance 1 to having variance $1+\sigma^{2}_{U}$, which would effectively mean we are in the standard measurement error setup, then the standard IV estimator would provide consistent estimates. [^19] We actually would not need to apply this transformation to whichever one is used as the instrument, since any affine transformation of the instrument will cancel out in the expression for the IV estimator. A convenient way to do this is to estimate $\sigma^{2}_{U}$ with $\nicefrac{{1}}{{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}-1]$, then note that since we want the variance of our predictor to be $1+\sigma^{2}_{U}$ we can multiply it by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ and use this transformed predictor in a standard IV analysis. [^20] Since the ratio of the uncorrected OLS and uncorrected IV estimators provides information about $\lambda$ (see [undefae]), yet another equivalent estimator can be constructed based on the square root of the product of the baseline OLS and uncorrected IV estimators using the partitioned $\hat{X}_{1}$ and $\hat{X}_{2}$. Since this approach is less intuitive, we do not discuss it here. Obviously, multiplying the predictor by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ has the result of dividing the estimated slope coefficient by the same factor, which is equivalent to multiplying it by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ as in Equation A.II.7 above.
As above, we may want to calculate our corrected IV estimator over all equally sized partitions of the indicators for estimating $\hat{X}_{1}$ and $\hat{X}_{2}$. For each such partition take the average of the corrected estimator using $\hat{X}_{1}$ to instrument for $\hat{X}_{2}$ and the one using $\hat{X}_{2}$ to instrument for $\hat{X}_{1}$, then choose as our estimator the median of all such partition averages. In fact, this estimator, which we call $\hat{\beta}^{*}_{\text{IV}}$, is equivalent to $\hat{\beta}_{\text{OLS}}^{*}$ (see Appendix II) In other words, the correlation-adjusted OLS estimator and the correlation-adjusted IV estimator are equivalent and below we refer to them together as the corrected estimator.
Despite the equivalence shown between $\hat{\beta}_{\text{OLS}}^{*}$ and $\hat{\beta}_{\text{IV}}^{*}$ in the bivariate case, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. If one predictor is measured with error, we may either have other “control” variables or the predictor measured with error may itself be a control variable in which case adjusting for this may be important to prevent bias in the other coefficient(s) of interest. [^21] Note in the latter case we may not care about appropriately adjusting the IV coefficient on the predictor that is measured with error. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this. In both the OLS and IV cases, we employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.
Finally, the framework here assumes classical (mean-zero, independent) measurement error. When non-classical error is a concern, identification can still sometimes be recovered under additional conditions; see, for example, [undefu]. In settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [undefaa]; these ideas complement our split-indicator strategy.
To see that $\hat{\beta}^{*}_{\text{IV}}$, is equivalent to $\hat{\beta}_{\text{OLS}}^{*}$ note that $\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}$, the linear regression slope coefficient estimate when predicting $\hat{X}_{1}$ with $\hat{X}_{2}$, is equal to $\text{Cor}(\hat{X}_{1},\hat{X}_{2})\sqrt{\frac{\text{V}(\hat{X}_{2})}{\text{V}(\hat{X}_{1})}}=\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ since the variance of both $\hat{X}_{1}$ and $\hat{X}_{2}$ are equal to 1 given that each is standardized. This also implies that $\hat{\beta}_{\text{OLS},\hat{X}_{2}\sim\hat{X}_{1}}=\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ since $\text{Cor}(\hat{X}_{1},\hat{X}_{2})=\text{Cor}(\hat{X}_{2},\hat{X}_{1})$ Therefore we can write the uncorrected IV slope coefficient using $\hat{X}_{1}$ as our predictor and $\hat{X}_{2}$ as our instrument as
$\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}}$ (A.II.8)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ (A.II.9)
and hence we have
$\hat{\beta}^{*}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (A.II.10)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}$
$=\hat{\beta}^{*}_{\text{OLS},\hat{X}_{2}}$
Appendix III: Additional Simulation Details & Results
A.III.1.1 Simulation Design Details
We adopt the same data-generating process as above; for each simulation iteration we:
  1. 1.
    Generate simulated matrix of indicators $W^{\textrm{Sim}}$ by randomly drawing from the predictive distribution of the ideal point model in Equation 8.
  2. 2.
    Estimate respondent ideal points $X$ based on the simulated indicators matrix $W^{\textrm{Sim}}$, rescaling these latent traits estimates to have mean 0 and variance 1 with higher values corresponding to more conservative positions, calling these estimates $X^{\textrm{Sim}}$.
  3. 3.
    Estimate the slope coefficient relating $X$ with $Y$ as specified in Equation 1 using simple OLS with the true latent predictor, with the estimated predictor, with the split sample IV (with uncorrected and two corrected estimators).
A.III.1.2 Ideology Estimation
Before discussing a larger-scale simulation that varies the number of observations and factors used in estimation, we first provide a simple illustration of the performance of the estimators discussed above through a set of simulations using a single fixed set of parameter values. To choose these “true” data-generating parameter values, we look at data from the 2016 Cooperative Congressional Election Survey (CCES). This survey includes a large number of policy questions on which respondents are asked their positions (e.g., whether respondents support background checks for all gun sales or whether they think women should always be allowed to have an abortion as a matter of choice). We use 30 of these policy questions as our indicators from which we estimate respondents’ ideologies (ideal points) and the item parameters (difficulty and discrimination parameters) based on the Clinton, Jackman, and Rivers (CJR) ideal point model. [^22] We use the binIRT function in the emIRT package. We drop several policy items that contain a large proportion of missing values. We also drop any respondents having missing values on any of these indicators or on the presidential vote question or who stated they did not vote or voted for someone other than Trump or Clinton, and then, to allow the simulations to run relatively quickly, we randomly select a subset of 1,000 remaining respondents to use. We then use the estimated ideal points to predict 2016 presidential vote (1=Trump, 0=Clinton).
We then set the true values for 1,000 respondents’ ideal points to those point estimates for $X_{i}$ (after imposing the identifying restriction that these ideal points have mean 0, variance 1, and that higher values correspond to more conservative positions) and set the true values of the difficulty ($\eta$) and discrimination parameters ($\alpha$) to those estimated from these indicators in our CCES sample. We also set the true values for the intercept $\beta_{0}$, the slope $\beta_{x}$, and the error variance $\textrm{V}(\epsilon_{i})$ to roughly correspond to the values estimated in the linear regression predicting presidential vote with the estimated ideal points using the baseline OLS approach. [^23] The estimated slope using our corrected estimates is slightly under 0.4, but for the purposes of the simulations, we use the value of 0.4 since a round number is easier to use as a baseline for comparison in figures and elsewhere. Note that this value for the slope is likely to be attenuated, i.e., smaller in magnitude than the true slope coefficient, since it was estimated from data that itself includes measurement error in the predictor, but we leave it as is for the true value in our simulations.
Although these simulations rely on a single set of true parameters, these values are based on real data of the sort that is often used in latent predictor applications. Furthermore, these simple simulations will illustrate some basic characteristics of the various estimators and their relative performance, at least in this specific example. Figure A.III.1 shows the distributions of simulated slope estimates ($\beta_{X}$) based on 100,000 simulation draws for each estimator.
The top pane of Figure A.III.1 provides a baseline, showing the simulated sampling distribution for the slope coefficient if we actually knew each respondent’s ideal point with certainty (i.e. if there was no measurement error in the predictor) and we estimated a simple bivariate linear regression predicting presidential vote with the ideal points $X$. As expected, the sampling distribution for the slope estimator is centered at the true slope of 0.4. The simulated RMSE is relatively small (0.01). Well over 60 percent of the simulations produced estimates within 0.01 of the true slope, and well over 90 percent produced estimates within 0.02 of the true slope. In a world where the true predictor is measured without error, we can obtain unbiased and relatively precise estimates of the slope coefficient of interest. Of course, with latent predictors, we will never know these true values.
The second pane in Figure A.III.1 plots the simulated distribution of the uncorrected OLS estimator in the case of measurement error. These simulations estimate the slope coefficient on $X$ predicting the presidential vote, assuming we do not actually observe the true latent trait $X$ but instead observe the set of 30 survey questions (indicators) from which we estimate $X$. In other words, we use estimated ideal points $\hat{X}$ as our predictor rather than the true ideal points without any correction or adjustment. As expected, the coefficient estimates are generally smaller in magnitude, with the average estimate falling just below 0.38. The RMSE is more than double that of the known $X$ case.
The third pane of Figure A.III.1 shows estimates from the standard split-sample IV approach without any corrections for the latent variable case. In these simulations, after simulating the matrix of indicators, we split the 30 indicators into the first 15 and the second 15, estimating ideal points for all respondents separately using each of these two sets. This produces $\hat{X_{1}}$ and $\hat{X_{2}}$, two separate measures. We then use two-stage least squares, using $X_{1}$ as an instrument for $X_{2}$ in the model predicting presidential vote. In the standard case of measurement error in an observable (as opposed to latent) predictor, this IV setup can yield consistent estimates of the slope coefficient. But as shown above, when applied to the latent case, it will over-correct for the attenuation bias, producing amplification bias. This is illustrated in the figure, where the average estimate is 0.43; the RMSE is nearly 4 times as large as in the known $X$ case (nearly double the value with uncorrected OLS and estimated $X$).
Finally, the bottom pane of Figure A.III.1 shows split-sample IV estimates corrected using the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ as described in Equation A.II.7. These estimates are quite similar to those using the true $X$, but with a very slightly higher bias and RMSE. This suggests the feasibility of estimating $\beta_{X}$ with the corrected IV approach in the more realistic case where we cannot use the true $X$ to estimate the measurement error variance of the split-half estimates.
In the last two applications (both corrected IV estimates), this simulation suggests favorable bias dynamics. (Appendix I discusses the possibility of a very enduring small amount of downward finite sample bias.)
This example illustrates the attenuation bias induced in estimates of a bivariate linear regression slope by using a noisy measure of a predictor. In our application, the measure is based on a relatively large number (30) of indicators, resulting in a relatively good signal-to-noise ratio. In cases where fewer indicators or less informative indicators are used, the attenuation bias will be larger, as seen below. It is also shown that while the standard IV approach to dealing with measurement error produces over-correction, it is possible to correct for this and obtain slope estimates that are nearly unbiased, with variance barely higher than if the true predictor were known without error.
Figure A.III.1: Illustrative simulation results. The true coefficient is highlighted in green (0.40). The mean of the estimates is highlighted in red across the various approaches.
A.III.1.3 Power Analysis
Figure A.III.2 shows that, across all the simulated settings examined, our corrected procedure retains close to nominal coverage while confidence intervals from the uncorrected OLS and IV estimators frequently fail to contain the true parameter value (0.4) even at moderate sample sizes. With the correct approach, as the number of indicators ($M$) grows, measurement error in the latent predictor diminishes, allowing the corrected estimator’s coverage to stay near nominal levels. In contrast, the uncorrected methods exhibit persistently lower coverage, particularly in smaller $M$ regimes. In this way, statistical inference is improved with the corrections described here.
Figure A.III.2: Power results varying NN and $M$.
A closer look at the small-$M$ regime reveals that even the corrected approach can show mildly reduced coverage there, reflecting finite-sample bias when estimating the latent factor from very little data. One possibility is that, in this low-information setting, random variation in the latent variable estimates via the bootstrap is insufficient to reflect true sampling uncertainty. Another possibility is the presence of small but lingering finite sample bias in this low-information case. Nonetheless, once $N$ and $M$ become sufficiently large, this small-$M$ artifact subsides; the corrected procedure once again not only outperforms the uncorrected OLS or IV, but also yields inferences approaching nominal levels.
A.III.1.4 Comparison with Additional Estimation Strategies
Researchers have increasingly turned to Bayesian strategies for inference regarding latent traits. We examine three such approaches: (1) method of composition, (2) multiple over-imputation, and (3) joint factor and outcome models. We compare these strategies using the same data-generating process outlined in the simulation section (see Section 5).
The method of composition approach draws from the posterior distribution of the latent trait estimates [undefao, undefam]. We first fit the outcome OLS model to each posterior draw of the latent traits. We then draw random samples from the resulting likelihood distribution, using draws from a Multivariate Gaussian centered at the point estimate and with a heteroskedasticity-consistent variance-covariance matrix. See [undefao] for details (especially the right panel of p. 215).
Multiple overimputation, described in [undefah] and refined in [undefh] for the joint outcome and measurement error modeling case, involves generating multiple plausible values for the latent traits using mean and variances from an initial MCMC step as priors—analyzing each imputed dataset separately before combining results.
In the joint modeling strategy, we fit a Bayesian model that simultaneously uses both outcome and observed factor information to generate estimates of the latent quantities of interest. This approach generates a Bayesian posterior over the coefficient estimates. Again, draws of the latent trait, $X$, are rescaled so that the posterior mean takes on a standard deviation of one.
Figure A.III.3: Various approaches, compared against the approaches analyzed in the main text.
Figure A.III.3 presents the results for these approaches. We observe that none of the Bayesian approaches outperform the sample splitting adjustment methods proposed in this paper. These results suggest that Bayesian modeling strategies do not necessarily solve the problems observed in this problem. The Bayesian posterior draws are noisier than the posterior mean estimate, hence increasing the degree of measurement error. Both the joint Bayesian model and multiple imputations face questions around how to perform standardization, whether that be within or across posterior draws.
In contrast, simple OLS estimates based on sample splitting, adjusted for measurement error as proposed in this paper, provide a compelling alternative. Beyond improvements in RMSE, this approach does not require fitting a larger-scale Bayesian model, thereby avoiding the increased computational costs associated with generating uncorrelated draws from the posterior distribution.
Appendix IV: Additional Empirical Results
A.IV.1.5 American Politics: Ideology & Presidential Vote
We here consider an application from American politics, in particular, from the literature on the relationship between latent ideology and partisan vote choice—a relationship that has been extensively studied and shown to be quite strong [undefv, undefan]. To what extent might measurement error in the estimation of latent ideology lead to over- or under-estimation in estimated relationships with the various strategies previously described?
We explore the relationship between latent ideology, as measured by 30 questions on the 2016 CCES regarding respondents’ policy positions, and vote choice for president. We again vary the number of observations used and policy factors considered in estimating latent traits, using random sub-samples and computing average estimates. Latent factors are estimated as described above using the expectation-maximization algorithm.
Figure A.IV.1: Measurement–error amplification for partisan‐ideal‐point predictors. Heat maps vary the number of observations (NN, xx‑axis) and the dimensionality of the latent‐space model ($M$, yy‑axis). Left panel: percentage increase (relative to naïve OLS) in the absolute magnitude of the vote‑share coefficient after applying the split‑sample correction; lighter shades indicate larger upward revisions. Right panel: split‑sample correlations between the two halves of the indicator set, with higher values signalling greater measurement stability.
Table A.IV.1: Percentage change in effect estimate of ideology on vote choice, corrected OLS vs. OLS. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 6 10 12 14 18 20 22 24 28 30
$n=$ 80 35.2 23.7 21.3 18.6 14.6 12.7 11.7 10.5 9.0 8.3
225 34.4 24.2 21.2 18.2 14.0 12.6 11.3 9.9 8.2 7.4
632 34.5 25.0 22.0 19.0 14.2 12.5 10.9 9.7 7.8 7.0
1778 34.6 25.4 22.2 19.1 14.1 12.4 10.9 9.5 7.7 6.9
5000 34.4 25.4 22.4 19.2 14.2 12.3 10.8 9.5 7.7 6.9
A.IV.1.6 Comparison: Corrected Estimator and Uncorrected IV in Democracy and Growth Application
We see in Table A.IV.2 that, if using the uncorrected IV strategy, we overestimate the effect of democracy on growth relative to the corrected IV strategy using a split sampling approach. This overestimation is largest when the number of observations and the number of democracy factors used in estimation are small. The underestimation decreases as the number of observations and the number of democracy factors increase, although again at a faster rate for the number of factors than for the number of observations. This again highlights the importance of gathering extensive information around the latent factor of interest.
Figure A.IV.2: Democracy–growth example: Sensitivity to the size of the indicator set. Left panel: percentage change in the estimated effect of democracy on long‑difference GDP growth after the split‑sample correction, plotted against the number of Freedom‑House sub‑indicators used ($M$). Right panel: mean split‑sample correlations for each value of $M$, showing how measurement error shrinks as more indicators are included.
Table A.IV.2: Percentage change in effect estimate of democracy on growth (long difference), corrected IV vs. uncorrected IV. Rows correspond to observation number; columns to dimensions used in estimating latent democratic status.
$M=$ 4 6 8 10 12 14 16
$n=$ 184 -4.8 -3.5 -2.6 -2.1 -1.7 -1.5 -1.3
A.IV.1.7 Effects of Political Knowledge
We here present additional empirical results on the application involving political knowledge.
Table A.IV.3: Percentage change in effect estimate of knowledge on turnout, corrected IV vs. OLS. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 4
$n=$ 80 49.6
225 50.7
632 48.7
1778 49.9
5000 49.8
Table A.IV.4: Percentage change in effect estimate of knowledge on turnout, corrected IV vs. uncorrected IV. Rows correspond to observation number; columns to dimensions used in estimating latent ideology.
$M=$ 4
$n=$ 80 -45.0
225 -41.3
632 -41.1
1778 -41.0
5000 -40.8
Figure A.IV.3: Left. Scatterplot of standardized sentiment scores computed from the first (x-axis) and second (y-axis) halves of each text. Points represent individual observations; the dashed diagonal line indicates perfect agreement; the solid line shows the linear regression fit with prediction intervals. Right. Coefficient by decade.
A.IV.1.8 Additional Information, Knowledge Example
We here provide the full text of questions from the American National Election Study (ANES) data used for the political knowledge and duty to vote application. These data were from the pre-release of the 2024 Time Series Study (downloaded 3/11/2025).
  • V241612
    For how many years is a United States Senator elected - that is, how many years are there in one full term of office for a U.S. Senator? [Respondents specify number of years, 6 is correct answer]
  • V241613
    On which of the following does the U.S. federal government currently spend the least? [Foreign aid (correct), Medicare, National defense, Social Security]
  • V241614
    Do you happen to know which party currently has the most members in the U.S. House of Representatives in Washington? [Democrats, Republicans (correct)]
  • V241615
    Do you happen to know which party currently has the most members in the U.S. Senate? [Democrats (correct), Republicans]
  • V241218x
    V241218x is a summary variable based on V241215, V241216 and V241217.
  • V241215
    Different people feel differently about voting. For some, voting is a duty - they feel they should vote in every election, regardless of their feelings about the candidates and parties. For others, voting is a choice - they feel free to vote or not to vote, depending on their feelings about the candidates and parties. For you personally, is voting mainly a duty, mainly a choice, or neither a duty nor a choice?[Mainly a duty, Mainly a choice, neither a duty nor a choice]
  • V241216
    How strongly do you feel that voting is a duty? [Very strongly, somewhat strongly, or not strongly / Not strongly, somewhat strongly, or very strongly]?
  • V241217
    How strongly do you feel that voting is a choice? [Very strongly, somewhat strongly, or not strongly / Not strongly, somewhat strongly, or very strongly]?
Appendix V: A lpmec Package Tutorial
lpmec is an R package for latent-predictor measurement error correction under identification restrictions (i.e., when a key regressor is an estimated latent trait from observed indicators, leading to attenuation bias in outcome regressions). Its core workflow repeatedly splits the indicator matrix into independent halves (and can average over many such partitions), estimates the latent score separately in each half (via EM IRT for binary/ordinal items using emIRT, PCA, simple averaging, or a user-supplied estimator), and then applies split-sample IV and OLS-style corrections that use the cross-half relationship to de-attenuate regression coefficients. The main entry points are lpmec_onerun() (single partition) and lpmec() (bootstrap + partition aggregation with percentile confidence intervals), with S3 print/summary/plot methods and built-in data (e.g., KnowledgeVoteDuty); the codebase also includes sensitivity reporting via sensemakr’s extreme robustness value.
For Bayesian estimation, lpmec supports MCMC either through an R backend (pscl::ideal) or an optional Python backend (numpyro via reticulate), plus an over-imputation variant using Amelia. The mcmc_joint option implements a joint Bayesian model that couples an IRT measurement model for the indicators with an outcome model (e.g., $Y_{i}\sim\mathcal{N}(\alpha+\beta\theta_{i},\sigma)$), so posterior uncertainty in the latent trait propagates directly into inference on the regression slope. When using the NumPyro/JAX backend, this can take advantage of JAX’s JIT/XLA compilation (and, optionally, GPU/TPU acceleration) and supports subsampling-based kernels for larger problems.
We average over n_partition split sample partitions:
# install.packages("devtools")
devtools::install_github(
  "cjerzak/lpmec-software",
  subdir = "lpmec"
library(lpmec)
data(KnowledgeVoteDuty)
# Latent error correction method, with partitioning and bootstrap
results <- lpmec::lpmec(
  Y = KnowledgeVoteDuty\$voteduty,
  observables = as.matrix(KnowledgeVoteDuty[, 2:5]),
  n_boot = 32L,
  n_partition = 10L,
  estimation_method = "em"
print(results)
summary(results)
Appendix VI: Code and Data Availability Statement
Additional data and materials will be made available online. For a code implementation, see
for open-source software implementing the methods described in the paper. Replication data and scripts will be made available at a Harvard Dataverse Repository.
Appendix VII: Measurement Error and Latent Predictors in the Political Science Literature
To assess the state of the political science literature regarding measurement error in latent predictors, we conducted an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP) as well as Political Analysis (PA) between 2000 and 2023. We observe in Figure A.VII.1 that the overall percentage of articles mentioning latent variables and, especially, measurement error is increasing, while the number of published works mentioning both measurement error and latent variables has remained relatively constant throughout the study period. This survey of recent political science articles demonstrates that while many researchers are aware of and concerned about measurement error, the full extent of its impact on analyses involving latent variables is, perhaps, not yet widely recognized or addressed in applied work.
Figure A.VII.1: While the overall number of mentions in top political science journals of measurement error and latent variables has been increasing, mentions of the intersection of the two have remained relatively constant in the survey period. The graph shows the number of articles containing each term by year for APSR, AJPS, JOP and PA from 2000 through 2023.